### v1

# A Survey on Multilingual Mental Disorders Detection from Social Media Data

Abstract
The increasing prevalence of mental health disorders globally highlights the urgent need for effective digital screening methods that can be used in multilingual contexts. Most existing studies, however, focus on English data, overlooking critical mental health signals that may be present in non-English texts. To address this important gap, we present the first survey on the detection of mental health disorders using multilingual social media data. We investigate the cultural nuances that influence online language patterns and self-disclosure behaviors, and how these factors can impact the performance of NLP tools. Additionally, we provide a comprehensive list of multilingual data collections that can be used for developing NLP models for mental health screening. Our findings can inform the design of effective multilingual mental health screening tools that can meet the needs of diverse populations, ultimately improving mental health outcomes on a global scale.
A Survey on Multilingual Mental Disorders Detection
from Social Media Data
Ana-Maria Bucur1,2,3, Marcos Zampieri4, Tharindu Ranasinghe5, Fabio Crestani3 1Interdisciplinary School of Doctoral Studies, University of Bucharest, Romania 2PRHLT Research Center, Universitat Politècnica de València, Spain 3Università della Svizzera italiana, Switzerland 4George Mason University, USA 5Aston University, Birmingham, UK
1 Introduction
It is estimated that nearly half of the population will develop at least one mental disorder by the age of 75 McGrath et al. (2023). Unfortunately, many people do not seek psychiatric help for mental health issues due to stigma, which manifests itself differently between cultures and is influenced by different cultural norms, religious beliefs and social attitudes Ahad et al. (2023). Due to the stigma associated with mental health and the limited access to professional care around the world, the World Health Organization (WHO) advocates for improved delivery of mental health services, including digital technologies to deliver remote care. [^1] https://www.who.int/news/item/17-06-2022-who-highlights-urgent-need-to-transform-mental-health-and-mental-health-care There is a pressing need for the integration of remote screening tools and the delivery of culturally adapted digital interventions Bond et al. (2023). Remote screening relies on processing language patterns associated with mental disorders, which can be identified from short essay writing Rude et al. (2004), text messages Nobles et al. (2018), or social media Eichstaedt et al. (2018).
The first well-known study on the detection of mental disorders using social media was conducted by De Choudhury et al. (2013). Multiple other studies have shown that the language used on Facebook can predict future depression diagnoses found in medical records, indicating that social media data could serve as a valuable complement to depression screening Eichstaedt et al. (2018). The current methods used for social media screening focus mainly on English data Skaik and Inkpen (2020); Harrigian et al. (2021). Additionally, there have been multiple workshops and shared tasks addressing NLP applications to mental health primarily on English data such as eRisk Parapar et al. (2024), CLPsych Chim et al. (2024) and LT-EDI Kayalvizhi et al. (2023).
There are important limitations in current NLP models when processing multilingual mental health-related data. Various studies analyzing English data from social media have shown that there are cultural differences in online language markers of mental disorders De Choudhury et al. (2017); Aguirre and Dredze (2021); Rai et al. (2024) and that the NLP models used for detection do not generalize on data from non-Western cultures Aguirre et al. (2021); Abdelkadir et al. (2024). Even one of the best predictors of depression in language, the use of the first person pronoun "I" Rude et al. (2004), for example, has different degrees of association with the severity of depression across different demographic groups Rai et al. (2024). This suggests that markers of mental disorders in social media language are not universal. One reason for this variation is that self-disclosure rates differ between cultures; collectivist cultures tend to have lower self-disclosure rates than individualist cultures in online settings Tokunaga (2009). Furthermore, non-native English speakers tend to use their native language for more intimate self-disclosures on social media, with higher rates of negative disclosure compared to posts in English Tang et al. (2011). This could have substantial implications for English-based social media screening tools, as they can overlook important signals of mental health disorders that are present in posts that are not written in English.
Recently, there have been efforts to develop detection models that focus on languages other than English, such as Portuguese Santos et al. (2024), German Zanwar et al. (2023), Arabic Almouzini et al. (2019), and Chinese Zhu et al. (2024). There have also been shared tasks specifically designed to address these issues, such as MentalRiskES Mármol-Romero et al. (2023), which focuses on the early detection of depression, suicide, and eating disorders in Spanish. To further contribute to these important efforts, we present the first survey on mental disorders detection from multilingual social media data. This survey aims to promote the development of multilingual NLP models that take into account cross-cultural and cross-language differences in online language.
This paper makes the following contributions:
  1. 1.
    We investigate cross-cultural and cross-language differences in the manifestations of mental disorders in social media.
  2. 2.
    We provide a comprehensive list of multilingual mental health datasets that capture linguistic diversity and can be used for developing multilingual NLP models. [^2] We make the list available and we will continuously update it: https://github.com/bucuram/multilingual-mental-health-datasets-nlp
  3. 3.
    We identify and describe several research gaps and future directions in the detection of multilingual mental disorders using online data.
2 Related Surveys
In this section, we analyze related surveys on the analysis of mental disorders from social media data. Calvo et al. (2017) is considered one of the first comprehensive surveys, presenting the datasets and NLP techniques used for mental health status detection and intervention. The survey explores research on various mental health conditions and states, including depression, mood disorders, psychological distress, and suicidal ideation, specifically in non-clinical texts such as user-generated content from social media and online forums. Similarly, recent surveys from Skaik and Inkpen (2020); Harrigian et al. (2021); Ríssola et al. (2021); Zhang et al. (2022); Garg (2023); Bucur et al. (2025) present the datasets, features, and models used to detect mental disorders from online content, focusing mainly on English language data. In addition to these surveys, Chancellor and De Choudhury (2020) provides a critical review of the study design and methods used to predict mental health status, along with recommendations to improve research in this field. Dhelim et al. (2023); Bucur et al. (2025) focus on studies that were published during the COVID-19 pandemic. It focuses on general mental well-being, loneliness, anxiety, stress, PTSD, depression, suicide, and other mental disorders.
Our paper fills an important gap in the literature by offering the first comprehensive survey of research on detecting mental disorders in languages other than English. The most related survey to ours is the one by Garg (2024), which focuses exclusively on low-resource languages. Our survey, however, has a broader scope as it discusses work on many languages irrespective of their resourcefulness.
3 Mental Disorders Detection Tasks Overview
Figure 1: Overview of tasks related to detecting mental health problems from social media.
In this section, we discuss the most common tasks related to predicting mental health disorders. When available, we include references to studies that focus on languages other than English. The prediction of mental health issues through social media is typically approached as a supervised classification task (Figure 1). The most common focus is on the binary classification of mental disorders. In this process, a collection of social media posts is used to train an NLP model, which then predicts a binary label that indicates the presence or absence of a mental disorder. Binary classification can be performed at the post-level, which is often used to predict suicidal ideation Huang et al. (2019) and depression Uddin et al. (2019). However, relying solely on a single post for decision making can lead to inaccurate predictions. Therefore, predictions can be made at the user level to detect conditions like depression Hiraga (2017), anxiety Zarate et al. (2023), bipolar disorder Sekulić et al. (2018), etc. Binary classification at the user level can also be modeled as an early risk prediction task, which aims to accurately label users as soon as possible, allowing the model to make a prediction or wait for more data before deciding Losada and Crestani (2016); Parapar et al. (2021).
Another important task is severity prediction, which can be modeled either as an ordinal regression / classification task or as a multiclass classification task. It is used primarily to predict the severity of depression Naseem et al. (2022); Kabir et al. (2023); Sampath and Durairaj (2022) or the risk of suicide attempts Benjachairat et al. (2024). Social media posts can be modeled longitudinally to detect moments of change in the mental health status of individuals. These shifts or escalations in mood can be used as a warning signal for potential suicidal behavior Tsakalidis et al. (2022b).
There are tasks designed to improve the explainability of the field, such as symptom prediction for mental disorders Liu et al. (2023); Yadav et al. (2020). Another step toward improving the explainability of model predictions is highlighting evidence for mental disorders Chim et al. (2024); Varadarajan et al. (2024). Mental health indicators from the social media timeline of an individual can be used to fill in validated questionnaires, with the goal of estimating symptoms of mental disorders that are usually assessed through survey-based methods such as the Beck’s Depression Inventory-II (BDI-II) [^3] https://naviauxlab.ucsd.edu/wp-content/uploads/2020/09/BDI21.pdf for depression assessment Parapar et al. (2021) or the Eating Disorder Examination Questionnaire (EDE-Q) [^4] https://www.corc.uk.net/media/1273/ede-q_quesionnaire.pdf for eating disorders Parapar et al. (2024).
Finally, mental health monitoring aggregated results from detection systems can be used to estimate the prevalence of mental disorders within a population. This approach was used during the COVID-19 pandemic to assess mental health burden, with results comparable to traditional survey-based methods Cohrdes et al. (2021).
4 Shared Tasks
Shared tasks have encouraged interdisciplinary collaborations between psychologists and computer scientists, resulting in systems that help detect mental disorders through social media analysis. These shared tasks have provided benchmark datasets that the research community continues to use, even beyond the official competitions.
The Early Detection of Mental Disorders Risk in Spanish (MentalRiskES) is the only shared task focused on detecting mental disorders in languages other than English. MentalRiskES includes tasks such as the detection of depression, anxiety, eating disorders, and suicidal risk in the Spanish language Mármol-Romero et al. (2023).
  Language Resource Datasets
  Arabic High Almouzini et al. (2019); Alghamdi et al. (2020); Alabdulkreem (2021); Musleh et al. (2022), CairoDep El-Ramly et al. (2021), Almars (2022); Maghraby and Ali (2022); Baghdadi et al. (2022), Arabic Dep 10,000 Helmy et al. (2024), Al-Haider et al. (2024); Abdulsalam et al. (2024); Al-Musallam and Al-Abdullatif (2022)
Chinese High Zhang et al. (2014); Huang et al. (2015); Cheng et al. (2017); Shen et al. (2018); Wu et al. (2018); Cao et al. (2019); Wang et al. (2019); Peng et al. (2019); Huang et al. (2019); Li et al. (2020), WU3D Wang et al. (2020), Yao et al. (2020); Yang et al. (2021); Chiu et al. (2021); Sun et al. (2022); Cai et al. (2023); Li et al. (2023); Guo et al. (2023); Wu et al. (2023); Lyu et al. (2023); Yu et al. (2023); Zhu et al. (2024)
French High Tabak and Purver (2020)
German High Cohrdes et al. (2021); Baskal et al. (2022); Tabak and Purver (2020), SMHD-GER Zanwar et al. (2023)
Japanese High Tsugawa et al. (2015); Hiraga (2017); Niimi (2021); Cha et al. (2022); Wang et al. (2023)
Spanish High Leis et al. (2019), SAD López-Úbeda et al. (2019), Valeriano et al. (2020); Ramírez-Cifuentes et al. (2020, 2021); Villa-Pérez et al. (2023), MentalRiskES Romero et al. (2024), Cremades et al. (2017); Coello-Guilarte et al. (2019)
Brazilian Portuguese Mid to High von Sperling and Ladeira (2019); Mann et al. (2020); Santos et al. (2020); de Carvalho et al. (2020), SetembroBR Santos et al. (2024), Mendes and Caseli (2024); Oliveira et al. (2024)
Dutch Mid to High Desmet and Hoste (2014, 2018)
Code-Mixed Hindi-English Mid to High Agarwal and Dhingra (2021)
Italian Mid to High Tabak and Purver (2020)
Korean Mid to High Lee et al. (2020); Park et al. (2020); Kim et al. (2022b, a); Cha et al. (2022)
Polish Mid to High Wołk et al. (2021)
Russian Mid to High Stankevich et al. (2019); Baskal et al. (2022); Narynov et al. (2020); Stankevich et al. (2020); Ignatiev et al. (2022)
Turkish Mid to High Baskal et al. (2022)
Bengali Mid Uddin et al. (2019); Victor et al. (2020); Kabir et al. (2022); Tasnim et al. (2022), BanglaSPD Islam et al. (2022), Ghosh et al. (2023); Hoque and Salma (2023), BSMDD Chowdhury et al. (2024)
Indonesian Mid Oyong et al. (2018); Yoshua and Maharani (2024)
Filipino Mid Tumaliuan et al. (2024); Astoveza et al. (2018)
Greek Mid Stamou et al. (2024)
Hebrew Mid Hacohen-Kerner et al. (2022)
Roman Urdu Mid Rehmani et al. (2024); Mohmand et al. (2024)
Thai Mid Katchapakirin et al. (2018); Hemtanon and Kittiphattanabawon (2019); Kumnunt and Sornil (2020); Hemtanon et al. (2020); Wongaptikaseree et al. (2020); Hämäläinen et al. (2021); Mahasiriakalayot et al. (2022); Boonyarat et al. (2024); Benjachairat et al. (2024)
Cantonese Low Gao et al. (2019)
Norwegian Low Uddin et al. (2022); Uddin (2022)
Sinhala Rare Rathnayake and Arachchige (2021), EmoMent Atapattu et al. (2022), Herath and Wijayasiriwardhane (2024)
Table 1: Available multilingual datasets for detecting mental disorders.
Other shared tasks are focused only on social media data in English. The Early Risk Prediction on the Internet Lab (eRisk) is an annual competition focusing mainly on the early detection of mental disorders, including depression, self-harm, pathological gambling, and eating disorders Parapar et al. (2024). The Workshop on Computational Linguistics and Clinical Psychology (CLPsych) includes various tasks, such as detecting depression and PTSD Coppersmith et al. (2015), labeling crisis posts Milne et al. (2016), and identifying moments of change Tsakalidis et al. (2022a) The Workshop on Language Technology for Equality, Diversity, and Inclusion (LT-EDI) organized tasks for predicting the severity of depression Kayalvizhi et al. (2023).
5 Methodology
To identify datasets for modeling the manifestations of mental disorders in languages other than English, we conducted a systematic search on major publication databases, including ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar. Initially, 405 studies were identified through database searches. After screening the abstracts, 215 papers were excluded because they did not mention the language of the data, or mention that the data is in English. Thus, following a review of the main body of the papers, the number of eligible studies was narrowed down to 108, which represents the final count of papers presenting datasets. Papers that did not present new data collections in languages other than English were excluded during the screening process. The PRISMA flow diagram for the survey is presented in Figure 3 in the Appendix.
6 Multilingual Datasets
The languages most frequently represented in the data collections are three high-resource languages: Chinese, Arabic, and Spanish. Although approximately half of the datasets were published in unranked venues, leading to low visibility for the research, the other half were published in high-ranking journals and conferences (Figure 4 in Appendix A.
6.1 Data Sources
Most of the datasets in English are sourced from Twitter [^5] All the datasets were collected before Twitter changed its name to X, so we refer to it as ‘Twitter’ in this paper. and Reddit Harrigian et al. (2021). Most non-English datasets in this section were also primarily collected from Twitter. However, Reddit was not as widely used for these data collections in non-English contexts. People use social media platforms differently. Twitter provides community and safety, helping raise awareness and combat stigma around mental health Berry et al. (2017). In contrast, Reddit allows for greater anonymity with “throwaway” accounts, encouraging users to openly share their experiences in detailed posts on specific subreddits De Choudhury and De (2014). This longer format supports post-level mental health analysis Chowdhury et al. (2024), while Twitter’s shorter posts favor user-level insights, requiring longitudinal data to identify patterns Tumaliuan et al. (2024). The data presented in this survey come from various populations and regions, and some of the sources are platforms that are exclusive to specific countries, such as Sina Weibo [^6] https://weibo.com used in China, VKontakte [^7] https://vk.com/ used in Russia, Pantip [^8] https://pantip.com/ in Thailand, or Everytime [^9] https://everytime.kr/ in Korea.
6.2 Languages
Table 1 presents all the datasets with multilingual data. A more detailed version of the table can be found in Appendix A, Table LABEL:tab:datasets. For classifying resource types, we used the framework proposed by Joshi et al. (2020). Figure 4 illustrates that most of the languages used in the data collections belong to some of the largest language families by number of speakers, specifically the Indo-European, Sino-Tibetan and Afro-Asiatic language families. The languages most frequently represented in the data collections are high-resource languages: Chinese appears in 25 data collections, Arabic is found in 11 datasets, and Spanish is included in 10 datasets. Even if most of the languages covered in the data are from high-, mid to high- and mid-resourced languages, we also have some languages with fewer resources, such as Cantonese and Norwegian. The Cantonese data collection was gathered by Gao et al. (2019) from Youtube comments and annotated for the risk of suicide. The Norwegian datasets related to depression were collected from a public online forum in Norway Uddin et al. (2022); Uddin (2022). Sinhala language, which was classified as rare by Joshi et al. (2020) is represented in three research papers. One of the papers contains Facebook data annotated for suicide ideation Herath and Wijayasiriwardhane (2024), while another contains depression-related data from Twitter and Facebook Rathnayake and Arachchige (2021). The third dataset contains data from Facebook, with more fine-grained labeled data on the presence of mental illness, anxiety, suicidal ideation, emotions, psychosomatic symptoms, and other manifestations Atapattu et al. (2022).
6.3 Mental Disorders
Figure 2 shows the distribution of mental disorders in different languages within the datasets. Depression is the most common mental disorder and is well-represented in the data. The languages that lack data on depression are Cantonese, Dutch, Hebrew, Hindi, and Turkish. Suicide is another mental disorder that frequently appears in collections. In contrast, the mental health problems that are least represented include eating disorders, obsessive-compulsive disorder (OCD), attention deficit / hyperactivity disorder (ADHD), autism spectrum disorder (ASD), anxiety, bipolar disorder, and schizophrenia.
Figure 2: Overview of the mental disorders addressed in each dataset, along with the annotation procedures.
6.4 Annotation Procedure
Most data collections were manually annotated (Figure 2). Manual annotation was carried out by mental health experts or psychologists Narynov et al. (2020); de Oliveira et al. (2022), graduate students who are native speakers of the language of interest Boonyarat et al. (2024); Uddin et al. (2019), or nonexpert individuals. However, some datasets do not specify who the annotators were or what guidelines they followed during the annotation process. Most datasets that collect user-level data from online platforms rely on the self-disclosure of mental health statuses. For example, they rely on explicit mentions of diagnoses (e.g. “I was diagnosed with depression") Tabak and Purver (2020); Villa-Pérez et al. (2023). The third most common annotation method involves asking social media users to complete validated questionnaires to diagnose mental disorders. The most frequently used survey-based methods include the CES-D Tsugawa et al. (2015); Lyu et al. (2023), BDI-II Sun et al. (2022); Stankevich et al. (2019); Ignatiev et al. (2022) or tools specifically designed for certain populations, such as the TMHQ [^10] Thai Mental Health Questionnaire Katchapakirin et al. (2018). Another reliable annotation approach is conducting clinical interviews to assess mental health problems Wołk et al. (2021). Less common and noisier annotation methods include identifying posts based on the presence of specific keywords López-Úbeda et al. (2019), by forum membership Agarwal and Dhingra (2021), or automatic annotation through another model trained on mental health data Cohrdes et al. (2021).
6.5 Availability of Data Collections
Of the 108 datasets listed in Table 1, only 23 are publicly available for download without any restrictions. These datasets focus on the detection of depression, suicide, and anorexia and are in various languages, including Arabic, Bengali, Brazilian Portuguese, Chinese, Hebrew, Hindi, Spanish, Russian, Roman Urdu, and Thai. For 15 of the datasets, access can be obtained by contacting the authors of the respective research papers, while four datasets require users to complete a data agreement to gain access. Additionally, four datasets are unavailable due to the sensitive nature of the data. For the remaining datasets, the research papers do not provide any information on data availability. Details about the availability of data collections can be found in Appendix A, Table LABEL:tab:datasets.
7 Mental Disorders Detection Approaches
In this section, we present the NLP methods proposed for the datasets in Section 6. Most approaches are monolingual and specifically target only one non-English language.
Classical approaches
Most approaches use Bag-of-Words, TF-IDF, or Word2Vec for text representation, which are then used as input for classical machine learning models Almouzini et al. (2019); Alghamdi et al. (2020); Helmy et al. (2024) or deep learning models Mann et al. (2020); Tasnim et al. (2022); Ghosh et al. (2023).
Pre-trained transformer-based models
While multilingual models like XLM-Roberta and Multilingual BERT demonstrate strong performance in downstream tasks, only two studies focus exclusively on these models Kabir et al. (2022); Hoque and Salma (2023). In contrast, twelve of the papers in Section 6 rely on pre-trained monolingual models specific to the target language, such as Chinese BERT Yao (2024), AraBERT Abdulsalam et al. (2024), German BERT Zanwar et al. (2023), Bangla BERT Chowdhury et al. (2024) and others. In addition, seven research papers evaluate both language-adapted and multilingual models Hacohen-Kerner et al. (2022); Oliveira et al. (2024).
Translation
Zahran et al. (2025) presented a comprehensive evaluation of LLMs on Arabic data related to depression, suicidal ideation, anxiety, and others. The authors found that LLMs performed better on original Arabic datasets compared to data that had been translated into English. Other works also rely on the detection using data translated from the target language to English Vajrobol et al. (2023). However, Schoene et al. (2025) has shown that automatically translating suicide dictionaries from English to low-resource languages often leads to spelling errors and fails to capture the cultural nuances of the speakers of the target language. When developing mental health models in other languages, some studies rely on translation from English to the target language, such as Greek Skianis et al. (2024) or various Indian languages Rajderkar and Bhat (2024).
Multilingual approaches
Methods developed for multiple languages simultaneously utilize cross-lingual embeddings and make use of information from languages with more mental health-related resources, such as English, to make predictions on Spanish data Coello-Guilarte et al. (2019). Lee et al. (2020) developed a cross-lingual model for suicidal ideation by translating data from Korean to English and Chinese. They used existing dictionaries related to suicidal ideation in these languages to inform predictions on the Korean language.
8 Cross-cultural and Cross-language Differences in Mental Health Expression
Culture influences the sources of distress, how it is expressed, how it is interpreted, the process of seeking help, and the responses of others Kirmayer et al. (2001). In addition, the way people perceive themselves influences their mental health. In Western cultures, there is a strong emphasis on personal narratives, and people tend to express their emotions more openly, a trend that is reflected in online posts Tokunaga (2009). In contrast, in Asian societies, individuals often internalize their emotional struggles or express them indirectly, influenced by their collectivist values Broczek et al. (2024). Although negative self-thoughts are a common characteristic of depression, in East Asian contexts, self-criticism is often viewed as a sign of healthy functioning Gotlib and Hammen (2008).
Symptoms of mental disorders
Cultural differences in the interpretation of mental health symptoms can lead individuals of certain backgrounds to minimize the psychological effects of mental distress. Instead, they may report more socially acceptable somatic symptoms Kirmayer et al. (2001). Somatic symptoms are common across various cultures, but the ways in which they are reported or understood can differ. In addition, there are culturally specific idioms of distress associated with mental disorders. One such example is the term “nervios” (translated as “nerves” in English), which is a syndrome of distress primarily studied in Latin American communities. This syndrome manifests with psychological and somatic symptoms and has a high comorbidity with anxiety and mood disorders De Snyder et al. (2000). The DSM-V American Psychiatric Association (2013), which is used for the assessment of mental disorders, includes cultural concepts of distress to help clinicians recognize how individuals from various cultures express psychological issues.
Mental health expressions in online language
Online expression varies between cultures and has been extensively studied among English-speaking individuals from different regions De Choudhury et al. (2017); Aguirre and Dredze (2021); Rai et al. (2024). When analyzing data from a peer-support mental health community, Loveys et al. (2018) found that manifestations of negative emotions differ between demographic groups. Moreover, Pendse et al. (2019) found that users in the US, UK, and Canada employed more clinical language to express mental distress compared to users from India, Malaysia, and the Philippines.
Variation of features across cultures
The tendency for self-focused attention, often referred to as “I”-language, is considered one of the strongest predictors of depression in language Mihalcea et al. (2024). As a result, the frequency of the pronoun “I” has been used in previous studies as a feature for detecting depression in English. However, it is crucial to carefully consider the applicability of this marker to non-English languages. This association has not been observed in non-Western individuals Rai et al. (2024) or in speakers of Chinese Lyu et al. (2023) or Romanian Trifu et al. (2024). While the pronoun “I” serves as a significant indicator of depression in English, its usage in other languages requires special attention due to linguistic differences. For example, English requires nouns or pronouns to be explicitly included as subjects in sentences. In contrast, some languages, such as Chinese and Romanian, are pro-drop languages, which allow the subject of the action to be omitted Koeneman and Zeijlstra (2019). This can result in a lower frequency of the personal pronoun “I” in these languages.
Mental health metaphors
Indicators of mental disorders are often displayed through metaphors. Depression is often described as weight, pressure, or darkness, and is often portrayed using containment metaphors Charteris-Black (2012). Metaphors are often used by individuals to articulate their experience and psychologists in the therapeutic process Mould et al. (2010). Mental illness metaphors have been extensively studied in English Charteris-Black (2012); Lazard et al. (2016) and have been used to predict mental states Shi et al. (2021); Zhang et al. (2021). With the exception of research in Spanish Coll-Florit and Climent (2023), there is a notable lack of resources to understand metaphors of mental illness in other languages.
It is essential to consider the various cultural and multilingual differences when developing automated methods to predict mental disorders based on language. These differences may explain why many studies have shown that models designed to predict mental illnesses often fail to generalize Aguirre et al. (2021); Abdelkadir et al. (2024).
9 Research Gaps
In this section, we highlight several research gaps that we hope will be explored in future studies.
Lack of mental health-related data for low-resource languages
As presented in Section 6, most data collection in non-English languages are often from mid- and high-resourced languages, with the exception of Cantonese, Norwegian, and Sinhala. Currently, many languages remain underrepresented, including high-resourced languages like French and mid-to-high resource languages such as Finnish, Croatian, and Vietnamese. Moreover, there is a lack of data collections for low-resource languages, which may hinder the development of online screening tools for individuals who speak these languages. Although few studies have used automatic translation for building datasets in languages other than English, it cannot accurately capture the cultural nuances of native speakers of the target language Schoene et al. (2025).
Cross-lingual expressions in underrepresented mental disorders
Although there are mental health-related datasets available in non-English data, most of them primarily focus on depression and suicide. Other mental disorders, such as anxiety, OCD, bipolar disorder, and PTSD, are underrepresented. To gain a better understanding of how these disorders manifest in the online language, the research community needs more linguistically diverse collections that encompass a wider range of mental disorders. This approach would not only facilitate a broader exploration of mental health expressions in various languages, but also help develop more inclusive and effective online mental health screening tools worldwide.
Multilingual approaches
As highlighted in Section 7, most NLP approaches have focused on processing data in a single target language, with multilingual approaches addressing multiple languages being almost nonexistent. Most existing NLP models developed for mental disorders detection do not support multiple languages effectively, which limits their applicability in multicultural and multilingual settings where mental health issues may manifest differently.
Annotation transparency and consistency
Although most of the datasets presented in this paper rely on manual annotation for labeling the data related to mental disorders, it is often unclear who did the annotations. The authors of the research papers should provide specific details about the annotation process, such as whether the annotators are mental health experts or non-experts, if they are native speakers of the target language, and whether they understand the cultural differences in the manifestations of mental disorders. These factors significantly impact the quality and reliability of the data, as understanding cultural nuances is essential in interpreting mental health expressions.
Explainability
While many mental health studies in English emphasize the importance of explainable approaches Yang et al. (2023a); Souto et al. (2023); Yang et al. (2023b), there is a significant opportunity for applying explainable approaches to non-English languages. Currently, few studies have examined model explainability in Bengali Ghosh et al. (2023) and Thai Vajrobol et al. (2023). These methods may help in understanding the various manifestations of mental disorders.
10 Conclusion
In this paper, we presented a comprehensive review of research for mental disorders detection from multilingual data sourced from social media. We highlight cross-cultural and multilingual differences in mental health expressions and provide a comprehensive list of data collections that can be used to develop multilingual NLP models for online mental health screening. Our focus was on non-English resources, as most previous research has focused on English Skaik and Inkpen (2020); Harrigian et al. (2021). Lastly, we identified several gaps in current research that we hope will be addressed in future interdisciplinary studies.
Future Directions and Call to Action
We aim to encourage researchers to develop mental health datasets in low-resource languages, fostering interdisciplinary collaborations with experts from psychology and mental health organizations, as seen in successful previous projects like REMO COST Action [^11] https://projects.tib.eu/remo, and PsyMine Ellendorff et al. (2016), which have primarily focused on English. By involving community members, multilingual shared tasks can be organized to identify mental disorders across different languages, inspired by successful SemEval multilingual tasks for offensive language Zampieri et al. (2020) and emotion detection Muhammad et al. (2025). Researchers can work together to annotate data in underrepresented languages while adhering to ethical protocols. By participating in these tasks, members of the ACL community can gain access to data collections that are essential for developing multilingual models. Such initiatives will improve the visibility of multilingual mental disorder detection and encourage further collaborations, providing researchers with more opportunities to address challenges in this field. Researchers can focus on building data collections for underrepresented mental disorders beyond depression and suicide, adhering to ethical guidelines and providing transparency in the annotation process Benton et al. (2017). Recent advances in explainability can also be applied to better understand the cultural manifestations of mental disorders.
Limitations
Our paper aims to provide a comprehensive review of cross-cultural language differences and the datasets available for developing multilingual NLP models. We included 108 data collections in this study and carefully reviewed each paper cited in our survey. However, it is possible that we may have overlooked some works that do not explicitly mention in their title or abstract that they focus on non-English languages.
Ethical Considerations
Data Collection
We recognize that using online data to identify mental disorders is a promising approach for early screening, but it also presents several ethical challenges Benton et al. (2017); Chancellor and De Choudhury (2020). To ensure that research protocols in this area comply with ethical guidelines, researchers must take the following steps: (1) obtain Institutional Review Board (IRB) approval, (2) follow ethical research protocols to protect sensitive data, as outlined by Benton et al. (2017), (3) obtain consent from participants, (4) de-anonymize the data and store it on a secure server. Any further sharing of the data with other researchers must adhere to the same ethical protocols. From our survey of 108 datasets, we found that only 18 received ethical approval from an IRB. In addition, 19 papers indicated that they anonymized the data to protect user privacy. It is concerning that only about 35% of the papers adhered to ethical practices in their research, highlighting the urgent need for a greater emphasis on ethical standards, especially since ethical disclosures were expected to gradually increase over time Ajmani et al. (2023).
Potential Consequences
Moreover, the ethical implications extend beyond data collection and storage. Researchers should consider the potential consequences of their findings on the populations studied and ensure that their work does not inadvertently stigmatize or harm individuals with mental health disorders. Incorrect predictions can have harmful effects on individuals’ lives. For instance, if a system falsely predicts that someone shows signs of mental disorders, it can adversely impact their well-being due to the stigma associated with such labels. This may lead individuals to believe there is something wrong with them, ultimately lowering their self-esteem Chancellor et al. (2019b). A false negative prediction occurs when the system fails to identify significant signs of distress, preventing the individual from receiving the necessary treatment or interventions. False negative predictions are particularly critical in cases of suicidal ideation, where a person’s life may be at risk. Chancellor et al. (2019a) critically discuss how subjects are represented in this area of research, highlighting the risk of inadvertently dehumanizing individuals. The language used in mental health-related papers can unintentionally perpetuate stigma, often referring to those involved in data collection as “sufferers” of mental disorders while labeling others as “normal.” Engaging with the community and stakeholders during the research process can help mitigate these risks and foster a more responsible approach to using online data in mental health research Chancellor et al. (2019b).
Model Validity
There are ongoing concerns regarding the construct validity of models trained on data collected from social media, specifically whether these models effectively measure the manifestations of mental disorders Chancellor and De Choudhury (2020). The datasets used in this survey predominantly rely on manual annotation or labeling through validated questionnaires, which are considered more reliable methods for annotation. However, it is essential to conduct interdisciplinary research and ground the constructs being measured in both theoretical and clinical frameworks. For example, clinical depression (or major depressive disorder) is fundamentally different from merely “feeling depressed.” The latter may refer to temporary feelings, while clinical depression encompasses a range of persistent symptoms. These symptoms may include depressed mood, loss of interest in previously enjoyed activities, changes in body weight, sleep disturbances, fatigue, psychomotor agitation or retardation, feelings of guilt, and thoughts of death or suicidal ideation American Psychiatric Association (2013). To be diagnosed with depression, these symptoms must be persistent and significantly impair an individual’s ability to function. Prioritizing interdisciplinary collaboration and rigorous validation methods is essential in addressing the complexities of mental health.
Representativeness
It is important to note that individuals active on social media represent only a subset of the overall population. As a result, there may be differences in how mental disorders are expressed among social media users compared to the general population. Using social media data can introduce bias, as it tends to reflect the experiences of younger and more technologically literate individuals who are more likely to engage with these platforms Chancellor et al. (2019b). In addition, datasets that include self-disclosure of a mental health diagnosis often come from individuals who are more likely to have sought professional help for their diagnosis and/or treatment. Furthermore, not everyone feels comfortable sharing sensitive information about their mental health online Chancellor et al. (2019b).
Cultural and Linguistic Variation
Understanding cultural and linguistic variations is crucial when developing automated methods for predicting mental disorders, as they help explain why many predictive models struggle to generalize effectively on data from different demographics Aguirre et al. (2021); Aguirre and Dredze (2021); Abdelkadir et al. (2024). Furthermore, each individual’s experience with depression is unique, and it is important to consider their distinct experiences and symptomatology. Algorithmic representations and abstractions play a crucial role in the understanding of mental illness and well-being by providing a framework for generalization Chancellor et al. (2019a). While these simplifications can help identify trends and better understand complex individual experiences, they also risk oversimplifying those experiences. It is important to recognize that generalizing can sometimes lead to misunderstandings regarding the unique nuances of mental health experiences and symptoms. Each person’s experience with mental health disorders is unique, and acknowledging this is essential for a deeper understanding of mental health.
Appendix A Appendix
A.1 Methodology details
Initially, 405 papers were retrieved through a database search across the ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar. After screening and assessing their eligibility, 108 papers were included in this survey. The PRISMA flow diagram is presented in Figure 3.
Figure 3: PRISMA flow diagram for our review.
A.2 Rankings of the publication venues for the multilingual datasets
Figure 4 presents an overview of these languages along with the ranking of the publications in which they appeared. The rankings for conferences are categorized as A, A, B, and C, following the CORE Rankings Portal. [^12] https://www.core.edu.au/conference-portal For journals, the rankings are classified as Q1, Q2, Q3, and Q4, based on the Journal Citation Reports. [^13] https://jcr.clarivate.com/ There are also datasets published in unranked conferences or journals. While about half of the datasets appeared in unranked venues, leading to lower visibility for the research, the other half were published in high-ranking journals and conferences.
Figure 4: Overview of the languages in the datasets, their language families, and the ranking of their publication venues.
Table 2: List of Non-English available datasets for mental disorders-related tasks using data posted on online platforms.
Dataset Language Mental disorder Platform Annotation Procedure Label Dataset Size Availab. Method Performance
Almouzini et al. (2019) Arabic depression Twitter Self-disclosure Binary 89 users, 2.7K posts UNK Bag-of-Unigrams, Linear SVM Accuracy: 87.5%, F1-score: 87.5%
Alghamdi et al. (2020) Arabic depression Online forums Manual annotation Binary 20K posts UNK Lexicon-based Accuracy: 80.45%, F1-score: 80.81%
Alabdulkreem (2021) Arabic depression Twitter Manual annotation Binary 200 users UNK Word2Vec, RNN-LSTM Accuracy: 72%, F1-score: 69%
Musleh et al. (2022) Arabic depression Twitter CES-D and self-disclosure Binary, DSM-5 symptoms 4.5K posts UNK TF-IDF, RF Accuracy: 82.39%, F1-score: 82.53%
CairoDep El-Ramly et al. (2021) Arabic depression Twitter, Reddit, Online forums Keywords, Manual annotation Binary 2.4K posts FREE AraBERT Accuracy: 96.93%, F1-score: 96.92%
Almars (2022) Arabic depression Twitter Manual annotation Binary 6.1K posts UNK Attention BiLSTM Accuracy: 83%, F1-score: 83%
Maghraby and Ali (2022) Arabic depression Twitter PHQ-9 PHQ-9 symptoms 1.2K posts FREE TF-IDF, RF F1-score: 98%
AraDepSu Hassib et al. (2022) Arabic depression, suicide Twitter Manual annotation Depression, depression with suicidal ideation, or non-depression 20K posts UNK MARBERT Accuracy: 91.20%, F1-score: 88.75%
Arabic Dep 10,000 Helmy et al. (2024) Arabic depression Twitter Manual annotation Binary 10K posts FREE TF-IDF, RBF SVM F1-score: 96.6%
Al-Haider et al. (2024) Arabic OCD Twitter Manual annotation Binary 8.7K posts UNK fastText, RF F1-score: 80%
Baghdadi et al. (2022) Arabic suicide Twitter Manual annotation Binary 2K posts FREE AraBERT Accuracy: 96.06%, F1-score: 95.86%
Abdulsalam et al. (2024) Arabic suicide Twitter Manual annotation Binary 5.7K posts UNK AraBERT Accuracy: 91%, F1-score: 88%
Al-Musallam and Al-Abdullatif (2022) Arabic depression Twitter Manual annotation Binary 6k posts UNK TF-IDF, LR Accuracy: 82%, F1-score: 81%
Uddin et al. (2019) Bengali depression Twitter Manual annotation Binary 1.1K posts FREE GRU Accuracy: 75.7%
Victor et al. (2020) Bengali depression Facebook, Twitter Manual annotation Binary 30K posts UNK TF-IDF, RF Accuracy: 90%
Kabir et al. (2022) Bengali depression Facebook Manual annotation Depression severity 5K posts FREE BiGRU Accuracy: 81%, F1-score: 81%
Tasnim et al. (2022) Bengali depression Facebook Manual annotation Binary 7K posts UNK BOW, TF-IDF, DT Accuracy: 97%, F1-score: 97%
BanglaSPD Islam et al. (2022) Bengali suicide Facebook Manual annotation Binary 1.7K posts UNK fastText, CNN-BiLSTM Accuracy: 61%, F1-score: 61%
Ghosh et al. (2023) Bengali depression Facebook, Twitter, YouTube Manual annotation Binary 15K posts AUTH fastText, BiLSTM-CNN Accuracy: 94.32%
Hoque and Salma (2023) Bengali depression Facebook Manual annotation Depression severity 2.5K posts UNK XLM-RoBERTa Accuracy: 61.11%, F1-score: 60.89%
BSMDD Chowdhury et al. (2024) Bengali depression Reddit, Twitter Manual annotation Binary 28K posts FREE GPT 3.5 Accuracy: 97.96%, F1-score: 98.04%
von Sperling and Ladeira (2019) Brazilian Portuguese depression Twitter Self-disclosure Binary 2.9K users UNK Hand-crafted features, SVM F1-score: 79.8%
Mann et al. (2020) Brazilian Portuguese depression Instagram BDI Binary 221 users UNK ELMo, ResNet, MLP F1-score: 79%
Santos et al. (2020) Brazilian Portuguese depression Twitter Self-disclosure Binary 224 users UNK TF-IDF, LR F1-score: 69%
de Carvalho et al. (2020) Brazilian Portuguese suicide Twitter Manual annotation Possibly/ Strongly concerning, Safe to ignore 2.4K posts UNK BERT-Portuguese F1-score: 79%
SetembroBR Santos et al. (2024) Brazilian Portuguese depression Twitter Self-disclosure Binary 18.8K users FREE BERTimbau F1-score: 63%
Mendes and Caseli (2024) Brazilian Portuguese depression symptoms Facebook Manual annotation Depression symptoms 780 posts UNK BERTimbau Precision: 76.14%
Oliveira et al. (2024) Brazilian Portuguese suicide Twitter Manual annotation Binary 3.7K posts FREE BERTimbau Accuracy: 96%
Gao et al. (2019) Cantonese suicide Youtube Manual annotation Binary 5K posts UNK Word2vec, LSTM Geometric mean of accuracies: 84.5%
Zhang et al. (2014) Chinese suicide Sina Weibo SPS SPS score 697 users UNK LIWC, LR RMSE: 11
Huang et al. (2015) Chinese suicide Sina Weibo Manual annotation Binary 7.3K posts UNK Topic modeling, LibSVM 80.0%
Cheng et al. (2017) Chinese suicide Sina Weibo Suicide Probability Scale (SPS), DASS-21 Binary 974 users UNK LIWC, SVM AUC: 0.61%
Shen et al. (2018) Chinese depression Sina Weibo Self-disclosure Binary 1.1K users UNK Hand-crafted features, DNN F1-score: 78.5%
Wu et al. (2018) Chinese depression Facebook CES-D Binary 1.4K users UNK Word2vec, Hand-crafted features ,RNN F1-score: 76.9%
Cao et al. (2019) Chinese suicide Sina Weibo Manual checking of self-report and/or appartenence to a suicide-related community Binary 7K users DUA fastText, RNN Accuracy: 91%
Wang et al. (2019) Chinese depression Sina Weibo Manual annotation Depression severity 13.9K users UNK BERT F1-score: 53.8%
Peng et al. (2019) Chinese depression Sina Weibo Manual annotation Binary 387 users UNK TF-IDF, SVM 83.46%
Huang et al. (2019) Chinese suicide Sina Weibo Manual annotation Binary 18.5K posts UNK LIWC, Dictionary, LR, DT, SVM F1-score: 0.88%
Li et al. (2020) Chinese depression Sina Weibo Self-disclosure Binary 1.8K users FREE Lexicon-based, RF F1-score: 76%
WU3D Wang et al. (2020) Chinese depression Sina Weibo Depression-related keywords Binary 32K users FREE XLNet mebeddings, BiGRU F1-score: 96.85%
Yao et al. (2020) Chinese depression Sina Weibo Manual, automatic annotation Binary 2.7K users UNK - -
Yang et al. (2021) Chinese depression Sina Weibo Manual annotation Depression severity 6.1K posts AUTH BERT-based F1-score: 65.7%
Chiu et al. (2021) Chinese, English depression Instagram Depression-related keywords Binary 520 users UNK Multimodal features, Adaboost F1-score: 83.5%
Sun et al. (2022) Chinese suicide, depression Sina Weibo BDI, SDS, Manual annotation Binary / Possibly/Strongly concerning, Safe to ignore 203 users, 1.2K posts UNK Gradient Boosting Accuracy: 82.4%
Cai et al. (2023) Chinese depression Sina Weibo Self-disclosure and manual annotation Binary 23K users FREE DNN F1-score: 92.02%
Li et al. (2023) Chinese depression Sina Weibo Self-disclosure, manual annotation Binary 4.8K users UNK Multimodal features, DNN F1-score: 92.78%
Guo et al. (2023) Chinese depression Sina Weibo Manual annotation Binary 3.1K users UNK Lexicon-based, XGBoost F1-score: 93.22%
Wu et al. (2023) Chinese suicide Dcard and PTT Manual annotation Risk levels 2K posts UNK - -
Lyu et al. (2023) Chinese depression Sina Weibo CES-D Binary 789 users AUTH LIWC, LR Pearson correlation: 0.33
Yu et al. (2023) Chinese anxiety Sina Weibo Self-Rating Anxiety Scale SAS score 1K users N/A LIWC, XGBoost Pearson correlation: 0.32
Zhu et al. (2024) Chinese anxiety Sina Weibo Manual annotation Binary 6K posts UNK LIWC, Word embeddings CNN F1-score: 86.13%
Wang et al. (2024) Chinese depression Sina Weibo Manual annotation Binary 14.8K users AUTH Multimodal features, DNN F1-score: 89.15%
Yao (2024) Chinese depression Sina Weibo Manual annotation Binary 200 users AUTH BERT, DNN Accuracy: 90%
Zhang et al. (2024) Chinese depression Sina Weibo Manual annotation Binary 1.6K users UNK Tencent Embeddings, HTN F1-score: 95.43%
Desmet and Hoste (2014) Dutch suicide Online forums Manual annotation Fine-grained labels 1.3K posts UNK BOW, SVM F1-score: 85.6%
Desmet and Hoste (2018) Dutch suicide Online forums Manual annotation Fine-grained labels 10K posts UNK BOW, Topic modeling, LibSVM F1-score: 92.69%
Abdelkadir et al. (2024) English, but from different populations depression Twitter Self-disclosure, Manual annotation Binary 531 users UNK MentalLongformer F1-score: 62%
Tumaliuan et al. (2024) Filipino, English depression Twitter PHQ-9 Binary 72 users AUTH - -
Astoveza et al. (2018) Filipino, Taglish suicide Twitter Manual annotation Binary 2.1K posts UNK BOW ,MLP Accuracy: 77.9%
Cohrdes et al. (2021) German depression Twitter Automatic annotation for PHQ-8 symptoms Binary 88K posts AUTH
SMHD-GER Zanwar et al. (2023) German depression, ADHD, anxiety, bipolar, OCD, PTSD, schizophrenia Reddit Manual annotation Labels for multiple disorders 28K posts DUA LIWC, BiLSTM F1-score: 52.22%
Baskal et al. (2022) German, Russian, Turkish, English eating disorders Reddit, Tumblr Manual annotation Binary 3K posts AUTH
Tabak and Purver (2020) German, French, Italian, Spanish, English depression Twitter Self-disclosure Binary 5K users UNK BOW, BiLSTM F1-score: 69%
Hacohen-Kerner et al. (2022) Hebrew anorexia Online forums Manual annotation Binary 200 posts FREE Hand-crafted features, RF Accuracy: 90.63%
Agarwal and Dhingra (2021) Code-Mixed Hindi-English suicide Reddit Subreddit membership Binary 6.4K posts FREE Indic BERT Accuracy: 98.54%
Oyong et al. (2018) Indonesian depression Twitter Manual annotation Binary 55 users UNK Hand-crafted depression score F1-score: 0.50%
Yoshua and Maharani (2024) Indonesian depression Twitter DASS-42 Binary 184 users UNK Word2Vec, DT F1-score: 94%
Tsugawa et al. (2015) Japanese depression Twitter CES-D, BDI Binary 209 users UNK Hand-crafted features, Topic modeling, SVM Accuracy: 66%
Hiraga (2017) Japanese depression Online blogs Self-disclosure Binary 101 users UNK Part-of-speech, NB Accuracy: 95.5%
Niimi (2021) Japanese depression TOBYO Blog theme Binary 901 users UNK TF-IDF, SVM F1-score: 96.2%
Wang et al. (2023) Japanese suicide Twitter Manual annotation Binary 30K posts N/A
Lee et al. (2020) Korean suicide Naver Cafe Membership in a forum Binary 31K posts UNK Word2Vec, RNN Accuracy: 87.49%
Park et al. (2020) Korean suicide Online forums Manual annotation Risk levels 2.7K posts AUTH XLM-R Accuracy: 88%
Kim et al. (2022a) Korean suicide Twitter Manual annotation Binary 20K posts, 414 users UNK
Kim et al. (2022b) Korean depression Online forums PHQ-9, Manual annotation PHQ-9 score, PHQ-9 symptoms 60 users, 28K posts UNK BERT-based Accuracy: 68.3%
Jung et al. (2023) Korean suicide Twitter Manual annotation Binary 20k posts UNK Metadata, word count, XGBoost F1-score: 83.57%
Cha et al. (2022) Korean, Japanese, English depression Twitter, Everytime Lexicon-based automatic annotation Binary 26M posts, 22K posts AUTH BERT-based F1-score:99%
Stamou et al. (2024) Modern Greek depression Twitter Self-disclosure Binary 78 users AUTH
Uddin (2022) Norwegian depression Online forums Manual annotation Binary 21.8K posts UNK TF-IDF, LSTM Accuracy: 99%
Uddin et al. (2022) Norwegian depression Online forums Manual annotation Binary 30K posts UNK Hand-crafted depression features; LSTM Accuracy: 99%
Wołk et al. (2021) Polish depression Facebook, Reddit Self-disclosure, clinical interview Binary 262 users UNK Hybrid Model; BERT Accuracy 71%
Rehmani et al. (2024) Roman Urdu depression Facebook Manual annotation Depression severity 3K posts AUTH SVM Accuracy: 84%
Mohmand et al. (2024) Roman Urdu depression Twitter Keywords-based annotations + Expert review Depression severity 25K posts FREE Transfer learning; BERT Accuracy: 99%
Stankevich et al. (2019) Russian depression VKontakte BDI BDI score 531 users UNK Psycholinguistic Markers F1 Score: 66%
Narynov et al. (2020) Russian depression VKontakte Manual annotation Binary 34K posts FREE
Stankevich et al. (2020) Russian depression VKontakte BDI BDI score 1.3K users UNK
Ignatiev et al. (2022) Russian depression VKontakte BDI Binary 619 users DUA CatBoost F1 Score: 69%
Rathnayake and Arachchige (2021) Sinhala depression Twitter, Facebook Manual annotation Binary 1K posts UNK KNN Accuracy: 70%
EmoMent Atapattu et al. (2022) Sinhala, English mental illness Facebook Manual annotation mental illness, sadness, suicidal, anxiety/stress, psychosomatic, other, irrelevant 2.8K posts AUTH RoBERTa F1 Score: 76%
Herath and Wijayasiriwardhane (2024) Sinhala suicide Facebook Manual annotation Binary 300 posts UNK Naive Bayes Accuracy: 79%
Leis et al. (2019) Spanish depression Twitter Self-disclosure, manual annotation Binary 540 users, 1K posts FREE
SAD López-Úbeda et al. (2019) Spanish anorexia Twitter Hashtags Binary 5.7K posts FREE SVM Accuracy: 91.6%
Valeriano et al. (2020) Spanish suicide Twitter Manual annotation Binary 2K posts FREE Word2Vec; LR Accuracy: 79%
Ramírez-Cifuentes et al. (2020) Spanish suicide Twitter Manual annotation Binary 252 users N/A
Ramírez-Cifuentes et al. (2021) Spanish anorexia Twitter Manual annotation Anorexia, control, under treatment, recovered, doubtful 645 users N/A
Villa-Pérez et al. (2023) Spanish, English depression, ADHD, anxiety, ASD, bipolar, eating disorders, OCD, PTSD, schizophrenia Twitter Self-disclosure Labels for multiple disorders 6K users DUA N-Grams; XGBoost AUC: 71.2%
MentalRiskES Romero et al. (2024) Spanish depression, anxiety, suicide, eating disorders Telegram Manual annotation Binary + suffer + in favour (sf), suffer + against (sa), suffer + other (so) for Depression 1.2K users AUTH Social media text; mDeBERTa F1 Score: 46%
Cremades et al. (2017) Spanish, English suicide Facebook, Twitter, Blogspot, Reddit, Pinterest Manual annotation Binary 97 posts FREE
Coello-Guilarte et al. (2019) Spanish, English depression Twitter Self-disclosure Binary 316 users FREE BA-LIWC F1 Score: 65%
Katchapakirin et al. (2018) Thai depression Facebook TMHQ Binary 35 users UNK RF F1 Score: 88.9%
Hemtanon and Kittiphattanabawon (2019) Thai depression Facebook Manual annotation Binary 1.5K posts UNK SVM F1 Score: 94%
Kumnunt and Sornil (2020) Thai depression Pantip Hashtags Binary 31K posts UNK CNN-LSTM F1 Score: 83.1%
Hemtanon et al. (2020) Thai depression Facebook PHQ-9 Binary 160 users UNK Social media features F1 Score: 91.4%
Wongaptikaseree et al. (2020) Thai depression Facebook TMHQ Binary 600 users UNK
Hämäläinen et al. (2021) Thai depression Online blogs Manual annotation Binary 900 posts FREE BERT Accuracy: 77.53%
Mahasiriakalayot et al. (2022) Thai depression Twitter Manual annotation Depression symptoms 3.1K posts UNK LSTM Accuracy: 89.17%
Boonyarat et al. (2024) Thai suicide Twitter Manual annotation Binary + 6 emotions 2.4K posts FREE Linguistic features; BERT F1-score: 90%
Benjachairat et al. (2024) Thai suicide Twitter Manual annotation C-SSRS Labels 5.1K posts AUTH Text features; LSTM F1-score: 93.88%

### v2

@@
 Abstract
-The increasing prevalence of mental health disorders globally highlights the urgent need for effective digital screening methods that can be used in multilingual contexts.
+The increasing prevalence of mental disorders globally highlights the urgent need for effective digital screening methods that can be used in multilingual contexts.
 Most existing studies, however, focus on English data, overlooking critical mental health signals that may be present in non-English texts.
-To address this important gap, we present the first survey on the detection of mental health disorders using multilingual social media data.
-We investigate the cultural nuances that influence online language patterns and self-disclosure behaviors, and how these factors can impact the performance of NLP tools.
-Additionally, we provide a comprehensive list of multilingual data collections that can be used for developing NLP models for mental health screening.
-Our findings can inform the design of effective multilingual mental health screening tools that can meet the needs of diverse populations, ultimately improving mental health outcomes on a global scale.
+To address this gap, we present a survey of the detection of mental disorders using social media data beyond the English language.
+We compile a comprehensive list of 108 datasets spanning 25 languages that can be used for developing NLP models for mental health screening.
+In addition, we discuss the cultural nuances that influence online language patterns and self-disclosure behaviors, and how these factors can impact the performance of NLP tools.
+Our survey highlights major challenges, including the scarcity of resources for low- and mid-resource languages and the dominance of depression-focused data over other disorders.
+By identifying these gaps, we advocate for interdisciplinary collaborations and the development of multilingual benchmarks to enhance mental health screening worldwide.
 A Survey on Multilingual Mental Disorders Detection
 from Social Media Data
-Ana-Maria Bucur1,2,3, Marcos Zampieri4, Tharindu Ranasinghe5, Fabio Crestani3 1Interdisciplinary School of Doctoral Studies, University of Bucharest, Romania 2PRHLT Research Center, Universitat Politècnica de València, Spain 3Università della Svizzera italiana, Switzerland 4George Mason University, USA 5Aston University, Birmingham, UK
+Ana-Maria Bucur1,2,3, Marcos Zampieri4, Tharindu Ranasinghe5, Fabio Crestani1 1Università della Svizzera italiana, Switzerland 2Interdisciplinary School of Doctoral Studies, University of Bucharest, Romania 3PRHLT Research Center, Universitat Politècnica de València, Spain 4George Mason University, USA 5Lancaster University, UK ana-maria.bucur-cosma@usi.ch
 1 Introduction
-It is estimated that nearly half of the population will develop at least one mental disorder by the age of 75 McGrath et al.
+Nearly half of the global population risks developing at least one mental disorder by age 75 McGrath et al.
 (2023).
-Unfortunately, many people do not seek psychiatric help for mental health issues due to stigma, which manifests itself differently between cultures and is influenced by different cultural norms, religious beliefs and social attitudes Ahad et al.
+Many individuals avoid seeking help due to stigma, which varies across cultures and is shaped by cultural norms, religious beliefs, and social attitudes Ahad et al.
 (2023).
-Due to the stigma associated with mental health and the limited access to professional care around the world, the World Health Organization (WHO) advocates for improved delivery of mental health services, including digital technologies to deliver remote care.
-[^1] https://www.who.int/news/item/17-06-2022-who-highlights-urgent-need-to-transform-mental-health-and-mental-health-care There is a pressing need for the integration of remote screening tools and the delivery of culturally adapted digital interventions Bond et al.
-(2023).
-Remote screening relies on processing language patterns associated with mental disorders, which can be identified from short essay writing Rude et al.
+Due to several challenges, including the stigma surrounding mental health, limited access in certain areas, economic factors, and a high global demand for services, the World Health Organization advocates for improved delivery of mental health services, including digital technologies to deliver remote care.
+[^1] https://www.who.int/news/item/17-06-2022-who-highlights-urgent-need-to-transform-mental-health-and-mental-health-care Furthermore, integrating remote screening tools and culturally adapted digital interventions is crucial Bond et al.
+(2023), as remote screening can detect language patterns linked to mental disorders from short essays Rude et al.
 (2004), text messages Nobles et al.
@@
 (2018).
+However, developing effective screening tools for mental disorders heavily relies on the availability of data.
+Therefore, in this work, we present the datasets that can be used to train models for screening in languages other than English.
 The first well-known study on the detection of mental disorders using social media was conducted by De Choudhury et al.
 (2013).
-Multiple other studies have shown that the language used on Facebook can predict future depression diagnoses found in medical records, indicating that social media data could serve as a valuable complement to depression screening Eichstaedt et al.
+Subsequent research has shown that the language used on Facebook can predict future depression diagnoses found in medical records, indicating that social media data could serve as a valuable complement to depression screening Eichstaedt et al.
 (2018).
-The current methods used for social media screening focus mainly on English data Skaik and Inkpen (2020); Harrigian et al.
-(2021).
-Additionally, there have been multiple workshops and shared tasks addressing NLP applications to mental health primarily on English data such as eRisk Parapar et al.
+Methods used for social media screening focus mainly on English data Skaik and Inkpen (2020); Harrigian et al.
+(2021), and workshops and shared tasks addressing NLP applications to mental health, such as eRisk Parapar et al.
+(2024); Crestani et al.
 (2022), CLPsych Chim et al.
 (2024) and LT-EDI Kayalvizhi et al.
-(2023).
-There are important limitations in current NLP models when processing multilingual mental health-related data.
-Various studies analyzing English data from social media have shown that there are cultural differences in online language markers of mental disorders De Choudhury et al.
+(2023) also primarily use English data.
+Current NLP models face major limitations in mental disorders detection in languages other than English.
+Studies show that cultural differences shape how mental disorders are expressed in language De Choudhury et al.
 (2017); Aguirre and Dredze (2021); Rai et al.
-(2024) and that the NLP models used for detection do not generalize on data from non-Western cultures Aguirre et al.
+(2024), which means that markers predictive in English often do not generalize well across different cultures Aguirre et al.
 (2021); Abdelkadir et al.
 (2024).
-Even one of the best predictors of depression in language, the use of the first person pronoun "I" Rude et al.
+Even one of the best predictors of depression in language, the use of the first person pronoun “I” Rude et al.
 (2004), for example, has different degrees of association with the severity of depression across different demographic groups Rai et al.
 (2024).
-This suggests that markers of mental disorders in social media language are not universal.
-One reason for this variation is that self-disclosure rates differ between cultures; collectivist cultures tend to have lower self-disclosure rates than individualist cultures in online settings Tokunaga (2009).
+This suggests that markers of mental disorders in language are not universal.
+In addition, self-disclosure rates vary across cultures; collectivist cultures tend to exhibit lower self-disclosure rates than individualist cultures in online settings Tokunaga (2009).
 Furthermore, non-native English speakers tend to use their native language for more intimate self-disclosures on social media, with higher rates of negative disclosure compared to posts in English Tang et al.
 (2011).
-This could have substantial implications for English-based social media screening tools, as they can overlook important signals of mental health disorders that are present in posts that are not written in English.
+This could have substantial implications for English-based social media screening tools, as they may overlook signals of mental disorders that are present in posts written in languages other than English.
 Recently, there have been efforts to develop detection models that focus on languages other than English, such as Portuguese Santos et al.
@@
 (2025), which focuses on the early detection of depression, suicide, and eating disorders in Spanish.
-To further contribute to these important efforts, we present the first survey on mental disorders detection from multilingual social media data.
-This survey aims to promote the development of multilingual NLP models that take into account cross-cultural and cross-language differences in online language.
+To further contribute to these important efforts, we present the first survey on the detection of mental disorders from social media data beyond English.
+This survey aims to promote the development of multilingual NLP models that account for cross-cultural and cross-linguistic differences in online language.
 This paper makes the following contributions:
   1. 1.
-    We investigate cross-cultural and cross-language differences in the manifestations of mental disorders in social media.
+    We provide a comprehensive list of multilingual mental health datasets that capture linguistic diversity and can be used for developing multilingual NLP models.
+[^2] We make the list publicly available, and we will continuously update it: https://github.com/bucuram/multilingual-mental-health-datasets-nlp
   2. 2.
-    We provide a comprehensive list of multilingual mental health datasets that capture linguistic diversity and can be used for developing multilingual NLP models.
-[^2] We make the list available and we will continuously update it: https://github.com/bucuram/multilingual-mental-health-datasets-nlp
+    We discuss cross-cultural and cross-language differences in the manifestations of mental disorders in social media.
   3. 3.
-    We identify and describe several research gaps and future directions in the detection of multilingual mental disorders using online data.
+    We identify and describe several research gaps and future directions in the detection of mental disorders using online data beyond English.
 2 Related Surveys
@@
 The survey explores research on various mental health conditions and states, including depression, mood disorders, psychological distress, and suicidal ideation, specifically in non-clinical texts such as user-generated content from social media and online forums.
-Similarly, recent surveys from Skaik and Inkpen (2020); Harrigian et al.
+Similarly, recent surveys Skaik and Inkpen (2020); Harrigian et al.
 (2021); Ríssola et al.
@@
 (2022); Garg (2023); Bucur et al.
-(2025) present the datasets, features, and models used to detect mental disorders from online content, focusing mainly on English language data.
+(2025a, b) present the datasets, features, and models used to detect mental disorders from online content, with a primary focus on English language data.
+Dhelim et al.
+(2023) and Bucur et al.
+(2025b) focus on studies that were published during the COVID-19 pandemic that address mental well-being, loneliness, anxiety, stress, PTSD, depression, suicide, and other mental disorders.
 In addition to these surveys, Chancellor and De Choudhury (2020) provides a critical review of the study design and methods used to predict mental health status, along with recommendations to improve research in this field.
-Dhelim et al.
+Our paper fills an important gap in the literature by offering the first comprehensive survey of research on detecting mental disorders in languages other than English.
+The most related survey to ours is the one conducted by Garg (2024), which focuses solely on low-resource languages and only discusses datasets in Thai, Bengali, Hindi, Japanese, and Korean.
+In contrast, our survey has a broader scope, covering research on a variety of languages, regardless of their resource availability.
+Our work complements previous surveys by providing an important data-centric foundation for researchers seeking to deploy and validate NLP methods in non-English contexts.
+3 Mental Disorders Detection Tasks Overview
+In this section, we discuss common tasks related to predicting mental disorders, emphasizing research available in non-English languages where applicable.
+The prediction of mental disorders via social media typically involves a supervised binary classification task, in which online posts are used to determine the presence or absence of disorders (Figure 1).
+This can be performed at the post level for predicting suicidal ideation Huang et al.
+(2019) and depression Uddin et al.
 (2019); Bucur et al.
-(2025) focus on studies that were published during the COVID-19 pandemic.
-It focuses on general mental well-being, loneliness, anxiety, stress, PTSD, depression, suicide, and other mental disorders.
-Our paper fills an important gap in the literature by offering the first comprehensive survey of research on detecting mental disorders in languages other than English.
-The most related survey to ours is the one by Garg (2024), which focuses exclusively on low-resource languages.
-Our survey, however, has a broader scope as it discusses work on many languages irrespective of their resourcefulness.
-3 Mental Disorders Detection Tasks Overview
-Figure 1: Overview of tasks related to detecting mental health problems from social media.
-In this section, we discuss the most common tasks related to predicting mental health disorders.
-When available, we include references to studies that focus on languages other than English.
-The prediction of mental health issues through social media is typically approached as a supervised classification task (Figure 1).
-The most common focus is on the binary classification of mental disorders.
-In this process, a collection of social media posts is used to train an NLP model, which then predicts a binary label that indicates the presence or absence of a mental disorder.
-Binary classification can be performed at the post-level, which is often used to predict suicidal ideation Huang et al.
-(2019) and depression Uddin et al.
+(2023), or at the user level for conditions like depression Hiraga (2017), anxiety Zarate et al.
+(2023), and bipolar disorder Sekulić et al.
 (2018).
-However, relying solely on a single post for decision making can lead to inaccurate predictions.
-Therefore, predictions can be made at the user level to detect conditions like depression Hiraga (2017), anxiety Zarate et al.
-(2023), bipolar disorder Sekulić et al.
-(2018), etc.
-Binary classification at the user level can also be modeled as an early risk prediction task, which aims to accurately label users as soon as possible, allowing the model to make a prediction or wait for more data before deciding Losada and Crestani (2016); Parapar et al.
+User-level prediction can serve as an early risk prediction task, facilitating timely assessments Losada and Crestani (2016); Bucur et al.
+(2021); Mármol-Romero et al.
 (2023).
-Another important task is severity prediction, which can be modeled either as an ordinal regression / classification task or as a multiclass classification task.
-It is used primarily to predict the severity of depression Naseem et al.
+Severity prediction assesses the intensity of conditions such as depression Naseem et al.
 (2022); Kabir et al.
-(2023); Sampath and Durairaj (2022) or the risk of suicide attempts Benjachairat et al.
+(2023); Sampath and Durairaj (2022) or suicide risk Benjachairat et al.
 (2024).
-Social media posts can be modeled longitudinally to detect moments of change in the mental health status of individuals.
-These shifts or escalations in mood can be used as a warning signal for potential suicidal behavior Tsakalidis et al.
-(2022b).
-There are tasks designed to improve the explainability of the field, such as symptom prediction for mental disorders Liu et al.
+Longitudinal analysis of social media helps detect moments of change in mental health Tsakalidis et al.
+(2022).
+Other tasks focus on symptom prediction Liu et al.
 (2023); Yadav et al.
-(2020).
-Another step toward improving the explainability of model predictions is highlighting evidence for mental disorders Chim et al.
+(2020) and highlighting evidence for mental health problems Chim et al.
 (2024); Varadarajan et al.
 (2024).
-Mental health indicators from the social media timeline of an individual can be used to fill in validated questionnaires, with the goal of estimating symptoms of mental disorders that are usually assessed through survey-based methods such as the Beck’s Depression Inventory-II (BDI-II) [^3] https://naviauxlab.ucsd.edu/wp-content/uploads/2020/09/BDI21.pdf for depression assessment Parapar et al.
-(2021) or the Eating Disorder Examination Questionnaire (EDE-Q) [^4] https://www.corc.uk.net/media/1273/ede-q_quesionnaire.pdf for eating disorders Parapar et al.
+Mental health indicators from the social media timeline of an individual can be used to fill in validated questionnaires, with the goal of estimating symptoms of mental disorders that are usually assessed through survey-based methods such as BDI-II [^3] https://naviauxlab.ucsd.edu/wp-content/uploads/2020/09/BDI21.pdf for depression assessment Parapar et al.
+(2021) or EDE-Q [^4] https://www.corc.uk.net/media/1273/ede-q_quesionnaire.pdf for eating disorders Parapar et al.
 (2024).
-Finally, mental health monitoring aggregated results from detection systems can be used to estimate the prevalence of mental disorders within a population.
-This approach was used during the COVID-19 pandemic to assess mental health burden, with results comparable to traditional survey-based methods Cohrdes et al.
+Lastly, mental health monitoring uses aggregated detection results to estimate the prevalence of disorders, as demonstrated during the COVID-19 pandemic Cohrdes et al.
 (2021).
-4 Shared Tasks
-Shared tasks have encouraged interdisciplinary collaborations between psychologists and computer scientists, resulting in systems that help detect mental disorders through social media analysis.
-These shared tasks have provided benchmark datasets that the research community continues to use, even beyond the official competitions.
-The Early Detection of Mental Disorders Risk in Spanish (MentalRiskES) is the only shared task focused on detecting mental disorders in languages other than English.
+Figure 1: Overview of tasks related to detecting mental health problems from social media.
+Shared tasks have encouraged interdisciplinary collaborations between psychologists and computer scientists.
+English-based shared tasks, such as eRisk Parapar et al.
+(2024), CLPsych Coppersmith et al.
+(2015), and LT-EDI Kayalvizhi et al.
+(2023) provided valuable benchmark datasets that the research community continues to use, even beyond the official competitions.
+MentalRiskES [^5] https://sites.google.com/view/mentalriskes2025 Mármol-Romero et al.
+(2025) is the only shared task focused on detecting mental disorders in languages other than English.
 MentalRiskES includes tasks such as the detection of depression, anxiety, eating disorders, and suicidal risk in the Spanish language Mármol-Romero et al.
 (2023).
+4 Methodology
+Figure 2: PRISMA flow diagram for our survey.
   Language Resource Datasets
@@
 Norwegian Low Uddin et al. (2022); Uddin (2022)
-Sinhala Rare Rathnayake and Arachchige (2021), EmoMent Atapattu et al. (2022), Herath and Wijayasiriwardhane (2024)
-Table 1: Available multilingual datasets for detecting mental disorders.
-Other shared tasks are focused only on social media data in English.
-The Early Risk Prediction on the Internet Lab (eRisk) is an annual competition focusing mainly on the early detection of mental disorders, including depression, self-harm, pathological gambling, and eating disorders Parapar et al.
-(2024).
-The Workshop on Computational Linguistics and Clinical Psychology (CLPsych) includes various tasks, such as detecting depression and PTSD Coppersmith et al.
-(2015), labeling crisis posts Milne et al.
-(2016), and identifying moments of change Tsakalidis et al.
-(2022a) The Workshop on Language Technology for Equality, Diversity, and Inclusion (LT-EDI) organized tasks for predicting the severity of depression Kayalvizhi et al.
-(2023).
-5 Methodology
-To identify datasets for modeling the manifestations of mental disorders in languages other than English, we conducted a systematic search on major publication databases, including ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar.
+Sinhala Low Rathnayake and Arachchige (2021), EmoMent Atapattu et al. (2022), Herath and Wijayasiriwardhane (2024)
+The classification of Sinhala is based on the work of Ranathunga and de Silva (2022), who identified the miscategorization of Sinhala in Joshi et al. (2020)’s classification.
+Table 1: Datasets for detecting mental disorders in languages other than English included in the current survey. The availability of resources is based on Joshi et al. (2020).
+Our primary objective is to catalog the available non-English resources derived from social media for the detection of mental disorders.
+We conducted a systematic search on major publication databases, including ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar.
 Initially, 405 studies were identified through database searches.
-After screening the abstracts, 215 papers were excluded because they did not mention the language of the data, or mention that the data is in English.
-Thus, following a review of the main body of the papers, the number of eligible studies was narrowed down to 108, which represents the final count of papers presenting datasets.
+After screening the abstracts, 215 papers were excluded because they did not mention the language of the data or mention that the data was in English.
+Thus, following a review of the main body of the papers, the number of eligible studies was narrowed down to 108, which represents the final count of papers included in this survey.
 Papers that did not present new data collections in languages other than English were excluded during the screening process.
-The PRISMA flow diagram for the survey is presented in Figure 3 in the Appendix.
+The PRISMA flow diagram is presented in Figure 2.
+The search terms used in our survey are provided in the Appendix A.
 5 Multilingual Datasets
-The languages most frequently represented in the data collections are three high-resource languages: Chinese, Arabic, and Spanish.
-Although approximately half of the datasets were published in unranked venues, leading to low visibility for the research, the other half were published in high-ranking journals and conferences (Figure 4 in Appendix A.
+In this section, we outline the datasets included in the current survey.
+The languages most frequently represented in these data collections are three high-resource languages: Chinese, Arabic, and Spanish.
+Although approximately half of the datasets were published in unranked venues, which diminishes their visibility, the other half were published in high-ranking journals and conferences (Figure 4 in Appendix A.
 5.1 Data Sources
@@
 However, Reddit was not as widely used for these data collections in non-English contexts.
-People use social media platforms differently.
+The data presented in this survey come from various populations and regions, and some of the sources are platforms that are exclusive to specific countries, such as Sina Weibo [^7] https://weibo.com used in China, VKontakte [^8] https://vk.com/ used in Russia, Pantip [^9] https://pantip.com/ in Thailand, or Everytime [^10] https://everytime.kr/ in Korea.
+Complete information about the sources of the data is presented in Table LABEL:tab:datasets in Appendix A.
+The way individuals use social media platforms to share information about themselves varies not only by platform but also by culture.
 Twitter provides community and safety, helping raise awareness and combat stigma around mental health Berry et al.
@@
 This longer format supports post-level mental health analysis Chowdhury et al.
-(2024), while Twitter’s shorter posts favor user-level insights, requiring longitudinal data to identify patterns Tumaliuan et al.
+(2024), while Twitter’s shorter posts favor user-level insights, requiring longitudinal data to identify language patterns Tumaliuan et al.
 (2024).
-The data presented in this survey come from various populations and regions, and some of the sources are platforms that are exclusive to specific countries, such as Sina Weibo [^6] https://weibo.com used in China, VKontakte [^7] https://vk.com/ used in Russia, Pantip [^8] https://pantip.com/ in Thailand, or Everytime [^9] https://everytime.kr/ in Korea.
+Moreover, even when individuals come from the same cultural background and speak the same language, the topics they choose to discuss are heavily influenced by the platform they use.
+For example, on Sina Weibo, users primarily talk about popular culture, while on Twitter, conversations in Chinese revolve more around politics Zhang and Gonçalves (2016).
+These differences, potentially shaped by platform governance and censorship, have direct implications for mental health data collection, as they influence the visibility, framing, and prevalence of mental health-related discourse across different platforms.
 5.2 Languages
-Table 1 presents all the datasets with multilingual data.
-A more detailed version of the table can be found in Appendix A, Table LABEL:tab:datasets.
-For classifying resource types, we used the framework proposed by Joshi et al.
-(2020).
-Figure 4 illustrates that most of the languages used in the data collections belong to some of the largest language families by number of speakers, specifically the Indo-European, Sino-Tibetan and Afro-Asiatic language families.
+Table 1 presents all the datasets included in this survey.
+To classify the availability of resource types, we used the framework proposed by Joshi et al.
+(2020), which categorizes resources into six classes (from Rare to High) based on the availability of both labeled and unlabeled NLP data.
 The languages most frequently represented in the data collections are high-resource languages: Chinese appears in 25 data collections, Arabic is found in 11 datasets, and Spanish is included in 10 datasets.
-Even if most of the languages covered in the data are from high-, mid to high- and mid-resourced languages, we also have some languages with fewer resources, such as Cantonese and Norwegian.
-The Cantonese data collection was gathered by Gao et al.
-(2019) from Youtube comments and annotated for the risk of suicide.
-The Norwegian datasets related to depression were collected from a public online forum in Norway Uddin et al.
-(2022); Uddin (2022).
-Sinhala language, which was classified as rare by Joshi et al.
-(2020) is represented in three research papers.
-One of the papers contains Facebook data annotated for suicide ideation Herath and Wijayasiriwardhane (2024), while another contains depression-related data from Twitter and Facebook Rathnayake and Arachchige (2021).
-The third dataset contains data from Facebook, with more fine-grained labeled data on the presence of mental illness, anxiety, suicidal ideation, emotions, psychosomatic symptoms, and other manifestations Atapattu et al.
-(2022).
+Even if most of the languages covered in the data are from high-, mid to high- and mid-resourced languages, we also have some languages with fewer resources, such as Cantonese, Norwegian and Sinhala.
+Most of the languages used in the data collections belong to some of the largest language families by number of speakers, specifically the Indo-European, Sino-Tibetan, and Afro-Asiatic language families.
+When analyzing the language resources based on speaker population and vitality according to Ethnologue Eberhard et al.
+(2023), we find that all the languages listed in Table 1 are categorized as large-institutional.
+This classification indicates that these languages are supported and maintained by institutions beyond their respective communities and have a large number of speakers.
+The categorization of all listed languages as large-institutional by Ethnologue underscores a robust support system that enhances their viability in NLP applications.
+However, it also highlights the urgent need to address disparities in resource availability for less-represented languages, promoting inclusivity in NLP research.
 5.3 Mental Disorders
@@
 In contrast, the mental health problems that are least represented include eating disorders, obsessive-compulsive disorder (OCD), attention deficit / hyperactivity disorder (ADHD), autism spectrum disorder (ASD), anxiety, bipolar disorder, and schizophrenia.
+Hazell et al.
+(2022)
 Figure 3: Overview of the mental disorders addressed in each dataset, along with the annotation procedures.
@@
 (2019); Ignatiev et al.
-(2022) or tools specifically designed for certain populations, such as the TMHQ [^10] Thai Mental Health Questionnaire Katchapakirin et al.
+(2022) or tools specifically designed for certain populations, such as the Thai Mental Health Questionnaire (TMHQ) Katchapakirin et al.
 (2018).
@@
 (2021).
+Manual annotation and annotation based on self-disclosure of diagnosis rely on individuals’ willingness to disclose personal information on social media.
+Research shows that collectivist cultures tend to have lower rates of self-disclosure compared to individualist cultures in online contexts Tokunaga (2009).
+However, there are notable differences within these cultures.
+For example, although Koreans are less likely to share personal information, they often provide more in-depth disclosures compared to individuals in the United States Yoo (2012).
+In addition, self-disclosure annotation relies on individuals explicitly naming their conditions, which may not occur in cultures where mental health is stigmatized.
+These annotation choices can have significant consequences for model behavior, potentially resulting in systematic false negatives for populations whose expressions of distress do not align with dominant annotation frameworks.
 5.5 Availability of Data Collections
-Of the 108 datasets listed in Table 1, only 23 are publicly available for download without any restrictions.
-These datasets focus on the detection of depression, suicide, and anorexia and are in various languages, including Arabic, Bengali, Brazilian Portuguese, Chinese, Hebrew, Hindi, Spanish, Russian, Roman Urdu, and Thai.
+Out of the 108 datasets listed in Table 1, only 23 are publicly available for download without restrictions.
+This is not surprising, as mental health datasets are often subject to privacy, ethical, and legal concerns, unlike many other NLP domains.
+These open datasets focus on the detection of depression, suicide, or anorexia in various languages, including Arabic, Bengali, Brazilian Portuguese, Chinese, Hebrew, Hindi, Spanish, Russian, Roman Urdu, and Thai.
 For 15 of the datasets, access can be obtained by contacting the authors of the respective research papers, while four datasets require users to complete a data agreement to gain access.
-Additionally, four datasets are unavailable due to the sensitive nature of the data.
+In addition, four datasets are unavailable due to the sensitive nature of the data.
 For the remaining datasets, the research papers do not provide any information on data availability.
@@
 6 Mental Disorders Detection Approaches
-In this section, we present the NLP methods proposed for the datasets in Section 6. Most approaches are monolingual and specifically target only one non-English language.
-Classical approaches
-Most approaches use Bag-of-Words, TF-IDF, or Word2Vec for text representation, which are then used as input for classical machine learning models Almouzini et al.
+In this section, we present the methods proposed for the datasets in Section 5.
+Most approaches are monolingual and specifically target only one non-English language.
+In Appendix A, Table LABEL:tab:datasets, we include detailed information on the methods used and evaluation scores for all the datasets.
+Classical Monolingual Approaches
+Most approaches rely on feature-based text representations such as Bag-of-Words, TF-IDF, and Word2Vec, combined with traditional classifiers like SVM, Logistic Regression Almouzini et al.
 (2019); Alghamdi et al.
@@
 (2023).
-Pre-trained transformer-based models
-While multilingual models like XLM-Roberta and Multilingual BERT demonstrate strong performance in downstream tasks, only two studies focus exclusively on these models Kabir et al.
-(2022); Hoque and Salma (2023).
-In contrast, twelve of the papers in Section 6 rely on pre-trained monolingual models specific to the target language, such as Chinese BERT Yao (2024), AraBERT Abdulsalam et al.
+These methods are straightforward and can be adapted to any language, making them suitable for the detection task.
+Monolingual pre-trained models
+With the rise of large pre-trained transformers, language-specific models such as Chinese BERT Yao (2024), AraBERT Abdulsalam et al.
 (2024), German BERT Zanwar et al.
 (2023), Bangla BERT Chowdhury et al.
-(2024) and others.
-In addition, seven research papers evaluate both language-adapted and multilingual models Hacohen-Kerner et al.
-(2022); Oliveira et al.
-(2024).
-Translation
-Zahran et al.
-(2025) presented a comprehensive evaluation of LLMs on Arabic data related to depression, suicidal ideation, anxiety, and others.
-The authors found that LLMs performed better on original Arabic datasets compared to data that had been translated into English.
-Other works also rely on the detection using data translated from the target language to English Vajrobol et al.
-(2023).
-However, Schoene et al.
-(2025) has shown that automatically translating suicide dictionaries from English to low-resource languages often leads to spelling errors and fails to capture the cultural nuances of the speakers of the target language.
-When developing mental health models in other languages, some studies rely on translation from English to the target language, such as Greek Skianis et al.
-(2024) or various Indian languages Rajderkar and Bhat (2024).
-Multilingual approaches
-Methods developed for multiple languages simultaneously utilize cross-lingual embeddings and make use of information from languages with more mental health-related resources, such as English, to make predictions on Spanish data Coello-Guilarte et al.
+(2024), BERTimbau Santos et al.
+(2024), IndicBERT Agarwal and Dhingra (2021) are being used for detecting mental disorders.
+However, a significant limitation of these pre-trained transformer-based models is that they are available only for a limited number of high- and mid-resource languages, rendering them impractical for use with low-resource languages.
+Multilingual models
+Methods developed for multiple languages simultaneously are rare.
+Some approaches use cross-lingual embeddings and leverage information from languages with more extensive mental health-related resources, such as English, to make predictions on Spanish data Coello-Guilarte et al.
 (2019).
@@
 (2020) developed a cross-lingual model for suicidal ideation by translating data from Korean to English and Chinese.
-They used existing dictionaries related to suicidal ideation in these languages to inform predictions on the Korean language.
+Although multilingual pre-trained models like XLM-Roberta and Multilingual BERT are frequently used, they often perform worse than language-specific models, such as BERTimbau and HeBERT Oliveira et al.
+(2024); Hacohen-Kerner et al.
+(2022).
+This discrepancy suggests that adapting models to specific cultural and linguistic contexts is often more effective than relying solely on general multilingual pre-training.
+Consequently, relying on a single model for multiple languages can increase the likelihood of errors.
+Translation-based
+Zahran et al.
+(2025) conducted a comprehensive evaluation of LLMs on Arabic data related to depression, suicidal ideation, anxiety, and other mental disorders.
+The authors showed that LLMs performed better on original Arabic datasets than on data that had been translated into English.
+Some studies also use data translated from the target language to English Vajrobol et al.
+(2023).
+When developing mental health models in languages other than English, certain studies use translations from English to the target language, as seen with Greek Skianis et al.
+(2024) and various Indian languages Rajderkar and Bhat (2024).
+However, Schoene et al.
+(2025) has shown that automatically translating suicide-related dictionaries from English to low-resource languages often results in spelling errors and fails to capture the cultural nuances of speakers in the target language.
+LLMs
+While LLMs have been successfully used in mental health tasks in English Wang et al.
+(2024b); Shah et al.
+(2025), research for other languages is limited.
+For instance, Zahran et al.
+(2025) evaluated LLMs on Arabic data and found that slight changes in instructions could harm performance, leading to parsing errors and prompt adherence issues.
+There are also concerns about increased hallucinations Lauscher et al.
+(2025) and unsafe responses Wang et al.
+(2024b) in languages other than English.
+Consequently, the use of LLMs in mental health tasks for non-English languages should be approached with caution.
+Performance across different languages
+As mentioned in Section 3, the detection of mental disorders is primarily approached as a classification task and evaluated using standard metrics such as Accuracy, Precision, Recall, and F1-score.
+Detailed performance scores can be found in Table LABEL:tab:datasets in Appendix A.
+The availability of resources has a significant impact on the performance of supervised or pre-trained language models.
+High-resource languages, such as Japanese, Chinese, and Arabic, generally yield strong results, indicating that larger and well-annotated datasets contribute to more reliable detection.
+In contrast, some languages, such as Sinhala, Brazilian Portuguese, and Bengali, show highly variable performance when comparing the detection of depression and suicidal ideation.
+For other languages included in this survey, such as French, German, Russian, and Italian, the performance in detecting depression is relatively low.
+Current methods demonstrate unequal effectiveness, which may leave speakers of certain languages without appropriate mental health screening tools.
+The success of these models is influenced by factors such as language-specific methodologies, the quality of datasets, the availability of resources, and cultural differences in expression.
+In addition, performance may vary depending on whether the task is binary classification, multi-class classification, or regression.
+Performance can also vary depending on whether the prediction is based on a single post (post-level classification) or multiple posts from the same user (user-level classification).
 7 Cross-cultural and Cross-language Differences in Mental Health Expression
@@
 In Western cultures, there is a strong emphasis on personal narratives, and people tend to express their emotions more openly, a trend that is reflected in online posts Tokunaga (2009).
-In contrast, in Asian societies, individuals often internalize their emotional struggles or express them indirectly, influenced by their collectivist values Broczek et al.
+In contrast, in Asian societies, individuals often internalize their emotional struggles or express them indirectly, influenced by their collectivist values Hu et al.
 (2014).
@@
 Symptoms of mental disorders
-Cultural differences in the interpretation of mental health symptoms can lead individuals of certain backgrounds to minimize the psychological effects of mental distress.
-Instead, they may report more socially acceptable somatic symptoms Kirmayer et al.
+Cultural differences in the interpretation of mental health symptoms can lead some individuals to minimize psychological distress, often reporting socially acceptable somatic symptoms such as body pain or fatigue Kirmayer et al.
 (2001).
-Somatic symptoms are common across various cultures, but the ways in which they are reported or understood can differ.
-In addition, there are culturally specific idioms of distress associated with mental disorders.
-One such example is the term “nervios” (translated as “nerves” in English), which is a syndrome of distress primarily studied in Latin American communities.
-This syndrome manifests with psychological and somatic symptoms and has a high comorbidity with anxiety and mood disorders De Snyder et al.
+While somatic symptoms are common across cultures, their expression and understanding can vary.
+Culturally specific idioms of distress, such as ”nervios” [^11] Translated as “nerves” in English.
+in Latin American communities, illustrate this, as they encompass both psychological and somatic symptoms and show a high comorbidity with anxiety and mood disorders De Snyder et al.
 (2000).
-The DSM-V American Psychiatric Association (2013), which is used for the assessment of mental disorders, includes cultural concepts of distress to help clinicians recognize how individuals from various cultures express psychological issues.
+The DSM-V APA (2013) includes cultural concepts of distress to aid clinicians in recognizing diverse expressions of psychological issues.
+In addition, Rai et al.
+(2025) found that Indian social media users on Reddit are more focused on seeking help than on discussing their symptoms or diagnoses.
 Mental health expressions in online language
@@
 (2024).
-While the pronoun “I” serves as a significant indicator of depression in English, its usage in other languages requires special attention due to linguistic differences.
+While the pronoun “I” serves as an indicator of depression in English, its usage in other languages requires special attention due to linguistic differences.
 For example, English requires nouns or pronouns to be explicitly included as subjects in sentences.
-In contrast, some languages, such as Chinese and Romanian, are pro-drop languages, which allow the subject of the action to be omitted Koeneman and Zeijlstra (2019).
-This can result in a lower frequency of the personal pronoun “I” in these languages.
+In contrast, some languages, such as Chinese and Romanian, are pro-drop languages, which allow the subject of the action to be omitted Koeneman and Zeijlstra (2019), resulting in a lower frequency of the personal pronoun “I” in these languages.
 Mental health metaphors
@@
 Depression is often described as weight, pressure, or darkness, and is often portrayed using containment metaphors Charteris-Black (2012).
-Metaphors are often used by individuals to articulate their experience and psychologists in the therapeutic process Mould et al.
+Metaphors are used by individuals to articulate their experiences, and psychologists use them in the therapeutic process Mould et al.
 (2010).
@@
 (2021).
-With the exception of research in Spanish Coll-Florit and Climent (2023), there is a notable lack of resources to understand metaphors of mental illness in other languages.
-It is essential to consider the various cultural and multilingual differences when developing automated methods to predict mental disorders based on language.
-These differences may explain why many studies have shown that models designed to predict mental illnesses often fail to generalize Aguirre et al.
+With the exception of research in Spanish Coll-Florit and Climent (2023), there is a lack of resources to understand metaphors of mental illness beyond English.
+Considering cultural and multilingual differences is crucial when developing automated methods for predicting mental disorders based on language.
+These variations may explain the failure of many predictive models to generalize effectively Aguirre et al.
 (2021); Abdelkadir et al.
@@
 In this section, we highlight several research gaps that we hope will be explored in future studies.
+Linguistic features used in English do not transfer reliably across languages
+Many commonly used linguistic markers of mental disorders, such as the use of first-person singular pronouns or sentiment polarity, are often considered universal features.
+However, these markers are actually dependent on language and shaped by cultural contexts, as previous research has indicated Trifan et al.
+(2020); Rai et al.
+(2024).
+While many studies have investigated markers of depression in language across cultures, most have focused primarily on individuals from various cultures who communicate in English Aguirre et al.
+(2021); Abdelkadir et al.
+(2024); Rai et al.
+(2025).
+There is a significant lack of research on how cross-cultural factors influence the expression of mental disorders in individuals’ native languages on social media, as well as how these factors affect the identification of such disorders.
 Lack of mental health-related data for low-resource languages
-As presented in Section 6, most data collection in non-English languages are often from mid- and high-resourced languages, with the exception of Cantonese, Norwegian, and Sinhala.
-Currently, many languages remain underrepresented, including high-resourced languages like French and mid-to-high resource languages such as Finnish, Croatian, and Vietnamese.
-Moreover, there is a lack of data collections for low-resource languages, which may hinder the development of online screening tools for individuals who speak these languages.
-Although few studies have used automatic translation for building datasets in languages other than English, it cannot accurately capture the cultural nuances of native speakers of the target language Schoene et al.
+As presented in Section 5, most datasets are often from mid- and high-resourced languages, with the exception of Cantonese, Norwegian, and Sinhala.
+Currently, many languages remain underrepresented, including high-resource languages like French and mid-to-high-resource languages such as Finnish, Croatian, and Vietnamese.
+Moreover, there is a lack of datasets in low-resource languages, which may hinder the development of online screening tools for individuals who speak these languages.
+While some studies have attempted to use automatic translation to build datasets in languages other than English, this approach often fails to accurately capture the cultural nuances of native speakers of the target language Schoene et al.
 (2025).
-Cross-lingual expressions in underrepresented mental disorders
-Although there are mental health-related datasets available in non-English data, most of them primarily focus on depression and suicide.
-Other mental disorders, such as anxiety, OCD, bipolar disorder, and PTSD, are underrepresented.
-To gain a better understanding of how these disorders manifest in the online language, the research community needs more linguistically diverse collections that encompass a wider range of mental disorders.
-This approach would not only facilitate a broader exploration of mental health expressions in various languages, but also help develop more inclusive and effective online mental health screening tools worldwide.
 Multilingual approaches
-As highlighted in Section 7, most NLP approaches have focused on processing data in a single target language, with multilingual approaches addressing multiple languages being almost nonexistent.
+As highlighted in Section 6, most NLP approaches have focused on processing data in a single target language, with multilingual approaches being almost nonexistent.
 Most existing NLP models developed for mental disorders detection do not support multiple languages effectively, which limits their applicability in multicultural and multilingual settings where mental health issues may manifest differently.
+While multilingual and cross-lingual approaches are often proposed as solutions to data scarcity, generalizability failures may stem not only from resource limitations but also from unmodeled cultural variation.
 Annotation transparency and consistency
-Although most of the datasets presented in this paper rely on manual annotation for labeling the data related to mental disorders, it is often unclear who did the annotations.
+Although most of the datasets presented in this paper rely on manual annotation for labeling the data related to mental disorders, it is often unclear who performed the annotations.
 The authors of the research papers should provide specific details about the annotation process, such as whether the annotators are mental health experts or non-experts, if they are native speakers of the target language, and whether they understand the cultural differences in the manifestations of mental disorders.
 These factors significantly impact the quality and reliability of the data, as understanding cultural nuances is essential in interpreting mental health expressions.
+For annotations based on questionnaires, it is crucial to use validated questionnaires specific to the target language rather than general English versions.
+Previous research has demonstrated that incorporating information about cultural idioms of distress in psychological assessments for mental disorders can enhance their validity Cork et al.
+(2019).
+Therefore, such considerations should also be integrated into the annotation process for mental disorders across cultures.
 Explainability
 While many mental health studies in English emphasize the importance of explainable approaches Yang et al.
-(2023a); Souto et al.
+(2023); Souto et al.
 (2023); Yang et al.
-(2023b), there is a significant opportunity for applying explainable approaches to non-English languages.
+(2024), there is a significant opportunity for applying explainable approaches to non-English languages.
+This could enhance our understanding of how mental disorders manifest culturally in social media language and provide insights into the diverse expressions of mental disorders.
 Currently, few studies have examined model explainability in Bengali Ghosh et al.
@@
 (2023).
-These methods may help in understanding the various manifestations of mental disorders.
 9 Conclusion
-In this paper, we presented a comprehensive review of research for mental disorders detection from multilingual data sourced from social media.
-We highlight cross-cultural and multilingual differences in mental health expressions and provide a comprehensive list of data collections that can be used to develop multilingual NLP models for online mental health screening.
+In this paper, we presented a comprehensive survey of research for mental disorders detection in social media data beyond English.
+We highlight cross-cultural and multilingual differences in mental health expressions and provide a comprehensive list of datasets that can be used to develop multilingual NLP models for online mental health screening.
 Our focus was on non-English resources, as most previous research has focused on English Skaik and Inkpen (2020); Harrigian et al.
@@
 Lastly, we identified several gaps in current research that we hope will be addressed in future interdisciplinary studies.
-Future Directions and Call to Action
-We aim to encourage researchers to develop mental health datasets in low-resource languages, fostering interdisciplinary collaborations with experts from psychology and mental health organizations, as seen in successful previous projects like REMO COST Action [^11] https://projects.tib.eu/remo, and PsyMine Ellendorff et al.
+Future Directions and Calls to Action
+We aim to encourage researchers to develop mental health datasets in languages other than English, fostering interdisciplinary collaborations with experts from psychology and mental health organizations, as seen in successful previous projects like REMO COST Action [^12] https://projects.tib.eu/remo, and PsyMine Ellendorff et al.
 (2016), which have primarily focused on English.
@@
 (2020) and emotion detection Muhammad et al.
-(2025).
+(2025a, b).
 Researchers can work together to annotate data in underrepresented languages while adhering to ethical protocols.
-By participating in these tasks, members of the ACL community can gain access to data collections that are essential for developing multilingual models.
+By participating in these tasks, members of the ACL community can gain access to datasets that are essential for developing multilingual models.
 Such initiatives will improve the visibility of multilingual mental disorder detection and encourage further collaborations, providing researchers with more opportunities to address challenges in this field.
-Researchers can focus on building data collections for underrepresented mental disorders beyond depression and suicide, adhering to ethical guidelines and providing transparency in the annotation process Benton et al.
+Researchers can focus on building datasets for underrepresented mental disorders beyond depression, adhering to ethical guidelines and providing transparency in the annotation process Benton et al.
 (2017).
@@
 Limitations
-Our paper aims to provide a comprehensive review of cross-cultural language differences and the datasets available for developing multilingual NLP models.
+Our paper aims to provide a comprehensive survey of cross-cultural language differences and the datasets available for developing multilingual NLP models.
 We included 108 data collections in this study and carefully reviewed each paper cited in our survey.
-However, it is possible that we may have overlooked some works that do not explicitly mention in their title or abstract that they focus on non-English languages.
+However, it is possible that we may have overlooked some works that do not explicitly mention in their title or abstract that they focus on languages other than English.
 Ethical Considerations
@@
 For example, clinical depression (or major depressive disorder) is fundamentally different from merely “feeling depressed.” The latter may refer to temporary feelings, while clinical depression encompasses a range of persistent symptoms.
-These symptoms may include depressed mood, loss of interest in previously enjoyed activities, changes in body weight, sleep disturbances, fatigue, psychomotor agitation or retardation, feelings of guilt, and thoughts of death or suicidal ideation American Psychiatric Association (2013).
+These symptoms may include depressed mood, loss of interest in previously enjoyed activities, changes in body weight, sleep disturbances, fatigue, psychomotor agitation or retardation, feelings of guilt, and thoughts of death or suicidal ideation APA (2013).
 To be diagnosed with depression, these symptoms must be persistent and significantly impair an individual’s ability to function.
@@
 Each person’s experience with mental health disorders is unique, and acknowledging this is essential for a deeper understanding of mental health.
+Acknowledgments
+We sincerely appreciate the insightful feedback provided by the anonymous reviewers, which has greatly contributed to the improvement of this paper.
 Appendix A Appendix
 A.1 Methodology details
-Initially, 405 papers were retrieved through a database search across the ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar.
-After screening and assessing their eligibility, 108 papers were included in this survey.
-The PRISMA flow diagram is presented in Figure 3.
-Figure 3: PRISMA flow diagram for our review.
+The search terms used in our survey are the following: (“mental disorder” OR “mental illness” ” OR “mental health” OR “depression” OR “anxiety” OR “bipolar” OR “ptsd” OR “post traumatic stress disorder” OR “suicide” OR “suicide ideation” OR “ocd” OR “obsessive compulsive disorder” OR “schizophrenia” OR “eating disorder” OR “anorexia” OR “bulimia”) AND “social media” AND (“multilingual” OR the list of languages provided in Joshi et al.
+(2020)).
 A.2 Rankings of the publication venues for the multilingual datasets
@@
 Almars (2022) Arabic depression Twitter Manual annotation Binary 6.1K posts UNK Attention BiLSTM Accuracy: 83%, F1-score: 83%
-Maghraby and Ali (2022) Arabic depression Twitter PHQ-9 PHQ-9 symptoms 1.2K posts FREE TF-IDF, RF F1-score: 98%
+Maghraby and Ali (2022) Arabic depression Twitter PHQ-9 PHQ-9 symptoms 1.2K posts FREE - -%
 AraDepSu Hassib et al.
-(2022) Arabic depression, suicide Twitter Manual annotation Depression, depression with suicidal ideation, or non-depression 20K posts UNK MARBERT Accuracy: 91.20%, F1-score: 88.75%
+(2022) Arabic depression, suicidal ideation Twitter Manual annotation Depression, depression with suicidal ideation, or non-depression 20K posts UNK MARBERT Accuracy: 91.20%, F1-score: 88.75%
 Arabic Dep 10,000 Helmy et al. (2024) Arabic depression Twitter Manual annotation Binary 10K posts FREE TF-IDF, RBF SVM F1-score: 96.6%
 Al-Haider et al. (2024) Arabic OCD Twitter Manual annotation Binary 8.7K posts UNK fastText, RF F1-score: 80%
-Baghdadi et al. (2022) Arabic suicide Twitter Manual annotation Binary 2K posts FREE AraBERT Accuracy: 96.06%, F1-score: 95.86%
-Abdulsalam et al. (2024) Arabic suicide Twitter Manual annotation Binary 5.7K posts UNK AraBERT Accuracy: 91%, F1-score: 88%
+Baghdadi et al. (2022) Arabic suicidal ideation Twitter Manual annotation Binary 2K posts FREE AraBERT Accuracy: 96.06%, F1-score: 95.86%
+Abdulsalam et al. (2024) Arabic suicidal ideation Twitter Manual annotation Binary 5.7K posts UNK AraBERT Accuracy: 91%, F1-score: 88%
 Al-Musallam and Al-Abdullatif (2022) Arabic depression Twitter Manual annotation Binary 6k posts UNK TF-IDF, LR Accuracy: 82%, F1-score: 81%
@@
 Tasnim et al. (2022) Bengali depression Facebook Manual annotation Binary 7K posts UNK BOW, TF-IDF, DT Accuracy: 97%, F1-score: 97%
-BanglaSPD Islam et al. (2022) Bengali suicide Facebook Manual annotation Binary 1.7K posts UNK fastText, CNN-BiLSTM Accuracy: 61%, F1-score: 61%
+BanglaSPD Islam et al. (2022) Bengali suicidal ideation Facebook Manual annotation Binary 1.7K posts UNK fastText, CNN-BiLSTM Accuracy: 61%, F1-score: 61%
 Ghosh et al. (2023) Bengali depression Facebook, Twitter, YouTube Manual annotation Binary 15K posts AUTH fastText, BiLSTM-CNN Accuracy: 94.32%
@@
 Santos et al. (2020) Brazilian Portuguese depression Twitter Self-disclosure Binary 224 users UNK TF-IDF, LR F1-score: 69%
-de Carvalho et al. (2020) Brazilian Portuguese suicide Twitter Manual annotation Possibly/ Strongly concerning, Safe to ignore 2.4K posts UNK BERT-Portuguese F1-score: 79%
+de Carvalho et al. (2020) Brazilian Portuguese suicidal ideation Twitter Manual annotation Possibly/ Strongly concerning, Safe to ignore 2.4K posts UNK BERT-Portuguese F1-score: 79%
 SetembroBR Santos et al. (2024) Brazilian Portuguese depression Twitter Self-disclosure Binary 18.8K users FREE BERTimbau F1-score: 63%
 Mendes and Caseli (2024) Brazilian Portuguese depression symptoms Facebook Manual annotation Depression symptoms 780 posts UNK BERTimbau Precision: 76.14%
-Oliveira et al. (2024) Brazilian Portuguese suicide Twitter Manual annotation Binary 3.7K posts FREE BERTimbau Accuracy: 96%
-Gao et al. (2019) Cantonese suicide Youtube Manual annotation Binary 5K posts UNK Word2vec, LSTM Geometric mean of accuracies: 84.5%
-Zhang et al. (2014) Chinese suicide Sina Weibo SPS SPS score 697 users UNK LIWC, LR RMSE: 11
-Huang et al. (2015) Chinese suicide Sina Weibo Manual annotation Binary 7.3K posts UNK Topic modeling, LibSVM 80.0%
-Cheng et al. (2017) Chinese suicide Sina Weibo Suicide Probability Scale (SPS), DASS-21 Binary 974 users UNK LIWC, SVM AUC: 0.61%
+Oliveira et al. (2024) Brazilian Portuguese suicidal ideation Twitter Manual annotation Binary 3.7K posts FREE GPT-4 F1-score: 98%
+Gao et al. (2019) Cantonese suicidal ideation Youtube Manual annotation Binary 5K posts UNK Word2vec, LSTM Geometric mean of accuracies: 84.5%
+Zhang et al. (2014) Chinese suicidal ideation Sina Weibo SPS SPS score 697 users UNK LIWC, LR RMSE: 11
+Huang et al. (2015) Chinese suicidal ideation Sina Weibo Manual annotation Binary 7.3K posts UNK Topic modeling, LibSVM F1-score: 80.0%
+Cheng et al. (2017) Chinese suicidal ideation Sina Weibo Suicide Probability Scale (SPS), DASS-21 Binary 974 users UNK LIWC, SVM AUC: 61%
 Shen et al. (2018) Chinese depression Sina Weibo Self-disclosure Binary 1.1K users UNK Hand-crafted features, DNN F1-score: 78.5%
 Wu et al. (2018) Chinese depression Facebook CES-D Binary 1.4K users UNK Word2vec, Hand-crafted features ,RNN F1-score: 76.9%
-Cao et al. (2019) Chinese suicide Sina Weibo Manual checking of self-report and/or appartenence to a suicide-related community Binary 7K users DUA fastText, RNN Accuracy: 91%
+Cao et al. (2019) Chinese suicidal ideation Sina Weibo Manual checking of self-report and/or appartenence to a suicide-related community Binary 7K users DUA fastText, RNN F1-score: 90.92%
 Wang et al. (2019) Chinese depression Sina Weibo Manual annotation Depression severity 13.9K users UNK BERT F1-score: 53.8%
-Peng et al. (2019) Chinese depression Sina Weibo Manual annotation Binary 387 users UNK TF-IDF, SVM 83.46%
-Huang et al. (2019) Chinese suicide Sina Weibo Manual annotation Binary 18.5K posts UNK LIWC, Dictionary, LR, DT, SVM F1-score: 0.88%
+Peng et al. (2019) Chinese depression Sina Weibo Manual annotation Binary 387 users UNK TF-IDF, SVM F1-score: 76.12%
+Huang et al. (2019) Chinese suicidal ideation Sina Weibo Manual annotation Binary 18.5K posts UNK LIWC, Dictionary, LR, DT, SVM F1-score: 88%
 Li et al. (2020) Chinese depression Sina Weibo Self-disclosure Binary 1.8K users FREE Lexicon-based, RF F1-score: 76%
@@
 Chiu et al. (2021) Chinese, English depression Instagram Depression-related keywords Binary 520 users UNK Multimodal features, Adaboost F1-score: 83.5%
-Sun et al. (2022) Chinese suicide, depression Sina Weibo BDI, SDS, Manual annotation Binary / Possibly/Strongly concerning, Safe to ignore 203 users, 1.2K posts UNK Gradient Boosting Accuracy: 82.4%
+Sun et al.
+(2022) Chinese suicidal ideation, depression Sina Weibo BDI, SDS, Manual annotation Binary / Possibly/Strongly concerning, Safe to ignore 203 users, 1.2K posts UNK Gradient Boosting F1-score: 78.90%
 Cai et al. (2023) Chinese depression Sina Weibo Self-disclosure and manual annotation Binary 23K users FREE DNN F1-score: 92.02%
@@
 Guo et al. (2023) Chinese depression Sina Weibo Manual annotation Binary 3.1K users UNK Lexicon-based, XGBoost F1-score: 93.22%
-Wu et al. (2023) Chinese suicide Dcard and PTT Manual annotation Risk levels 2K posts UNK - -
+Wu et al. (2023) Chinese suicidal ideation Dcard and PTT Manual annotation Risk levels 2K posts UNK - -
 Lyu et al. (2023) Chinese depression Sina Weibo CES-D Binary 789 users AUTH LIWC, LR Pearson correlation: 0.33
@@
 Zhu et al. (2024) Chinese anxiety Sina Weibo Manual annotation Binary 6K posts UNK LIWC, Word embeddings CNN F1-score: 86.13%
-Wang et al. (2024) Chinese depression Sina Weibo Manual annotation Binary 14.8K users AUTH Multimodal features, DNN F1-score: 89.15%
+Wang et al. (2024a) Chinese depression Sina Weibo Manual annotation Binary 14.8K users AUTH Multimodal features, DNN F1-score: 89.15%
 Yao (2024) Chinese depression Sina Weibo Manual annotation Binary 200 users AUTH BERT, DNN Accuracy: 90%
 Zhang et al. (2024) Chinese depression Sina Weibo Manual annotation Binary 1.6K users UNK Tencent Embeddings, HTN F1-score: 95.43%
-Desmet and Hoste (2014) Dutch suicide Online forums Manual annotation Fine-grained labels 1.3K posts UNK BOW, SVM F1-score: 85.6%
-Desmet and Hoste (2018) Dutch suicide Online forums Manual annotation Fine-grained labels 10K posts UNK BOW, Topic modeling, LibSVM F1-score: 92.69%
+Desmet and Hoste (2014) Dutch suicidal ideation Online forums Manual annotation Fine-grained labels 1.3K posts UNK BOW, SVM F1-score: 85.6%
+Desmet and Hoste (2018) Dutch suicidal ideation Online forums Manual annotation Fine-grained labels 10K posts UNK BOW, Topic modeling, LibSVM F1-score: 92.69%
 Abdelkadir et al. (2024) English, but from different populations depression Twitter Self-disclosure, Manual annotation Binary 531 users UNK Mental Longformer F1-score: 62%
 Tumaliuan et al. (2024) Filipino, English depression Twitter PHQ-9 Binary 72 users AUTH - -
-Astoveza et al. (2018) Filipino, Taglish suicide Twitter Manual annotation Binary 2.1K posts UNK BOW ,MLP Accuracy: 77.9%
+Astoveza et al. (2018) Filipino, Taglish suicidal ideation Twitter Manual annotation Binary 2.1K posts UNK BOW, MLP Accuracy: 77.9%
 Cohrdes et al. (2021) German depression Twitter Automatic annotation for PHQ-8 symptoms Binary 88K posts AUTH
-SMHD-GER Zanwar et al. (2023) German depression, ADHD, anxiety, bipolar, OCD, PTSD, schizophrenia Reddit Manual annotation Labels for multiple disorders 28K posts DUA LIWC, BiLSTM F1-score: 52.22%
+SMHD-GER Zanwar et al. (2023) German depression, ADHD, anxiety, bipolar, OCD, PTSD, schizophrenia Reddit Manual annotation Labels for multiple disorders 28K posts DUA LIWC, BiLSTM F1-score: 50.89%
 Baskal et al. (2022) German, Russian, Turkish, English eating disorders Reddit, Tumblr Manual annotation Binary 3K posts AUTH
@@
 Hacohen-Kerner et al. (2022) Hebrew anorexia Online forums Manual annotation Binary 200 posts FREE Hand-crafted features, RF Accuracy: 90.63%
-Agarwal and Dhingra (2021) Code-Mixed Hindi-English suicide Reddit Subreddit membership Binary 6.4K posts FREE Indic BERT Accuracy: 98.54%
-Oyong et al. (2018) Indonesian depression Twitter Manual annotation Binary 55 users UNK Hand-crafted depression score F1-score: 0.50%
+Agarwal and Dhingra (2021) Code-Mixed Hindi-English suicidal ideation Reddit Subreddit membership Binary 6.4K posts FREE Indic BERT Accuracy: 98.54%
+Oyong et al. (2018) Indonesian depression Twitter Manual annotation Binary 55 users UNK Hand-crafted depression score F1-score: 50%
 Yoshua and Maharani (2024) Indonesian depression Twitter DASS-42 Binary 184 users UNK Word2Vec, DT F1-score: 94%
@@
 Niimi (2021) Japanese depression TOBYO Blog theme Binary 901 users UNK TF-IDF, SVM F1-score: 96.2%
-Wang et al. (2023) Japanese suicide Twitter Manual annotation Binary 30K posts N/A
-Lee et al. (2020) Korean suicide Naver Cafe Membership in a forum Binary 31K posts UNK Word2Vec, RNN Accuracy: 87.49%
-Park et al. (2020) Korean suicide Online forums Manual annotation Risk levels 2.7K posts AUTH XLM-R Accuracy: 88%
-Kim et al. (2022a) Korean suicide Twitter Manual annotation Binary 20K posts, 414 users UNK
-Kim et al. (2022b) Korean depression Online forums PHQ-9, Manual annotation PHQ-9 score, PHQ-9 symptoms 60 users, 28K posts UNK BERT-based Accuracy: 68.3%
-Jung et al. (2023) Korean suicide Twitter Manual annotation Binary 20k posts UNK Metadata, word count, XGBoost F1-score: 83.57%
+Wang et al. (2023) Japanese suicidal ideation Twitter Manual annotation Binary 30K posts N/A
+Lee et al. (2020) Korean suicidal ideation Naver Cafe Membership in a forum Binary 31K posts UNK Word2Vec, RNN F1-score: 87.49%
+Park et al. (2020) Korean suicidal ideation Online forums Manual annotation Risk levels 2.7K posts AUTH KoBERT F1-score: 88%
+Kim et al. (2022a) Korean suicidal ideation Twitter Manual annotation Binary 20K posts, 414 users UNK
+Kim et al. (2022b) Korean depression Online forums PHQ-9, Manual annotation PHQ-9 score, PHQ-9 symptoms 60 users, 28K posts UNK BERT-based F1-score: 93%
+Jung et al. (2023) Korean suicidal ideation Twitter Manual annotation Binary 20k posts UNK Metadata, word count, XGBoost F1-score: 83.57%
 Cha et al. (2022) Korean, Japanese, English depression Twitter, Everytime Lexicon-based automatic annotation Binary 26M posts, 22K posts AUTH BERT-based F1-score: 99%
 Stamou et al. (2024) Modern Greek depression Twitter Self-disclosure Binary 78 users AUTH
-Uddin (2022) Norwegian depression Online forums Manual annotation Binary 21.8K posts UNK TF-IDF, LSTM Accuracy: 99%
-Uddin et al. (2022) Norwegian depression Online forums Manual annotation Binary 30K posts UNK Hand-crafted depression features; LSTM Accuracy: 99%
+Uddin (2022) Norwegian depression Online forums Manual annotation Binary 21.8K posts UNK TF-IDF, LSTM F1-score: 97%
+Uddin et al. (2022) Norwegian depression Online forums Manual annotation Binary 30K posts UNK Hand-crafted depression features; LSTM F1-score: 98%
 Wołk et al. (2021) Polish depression Facebook, Reddit Self-disclosure, clinical interview Binary 262 users UNK Hybrid Model; BERT Accuracy 71%
-Rehmani et al. (2024) Roman Urdu depression Facebook Manual annotation Depression severity 3K posts AUTH SVM Accuracy: 84%
-Mohmand et al. (2024) Roman Urdu depression Twitter Keywords-based annotations + Expert review Depression severity 25K posts FREE Transfer learning; BERT Accuracy: 99%
-Stankevich et al. (2019) Russian depression VKontakte BDI BDI score 531 users UNK Psycholinguistic Markers F1 Score: 66%
+Rehmani et al. (2024) Roman Urdu depression Facebook Manual annotation Depression severity 3K posts AUTH SVM F1-score: 70.3%
+Mohmand et al. (2024) Roman Urdu depression Twitter Keywords-based annotations + Expert review Depression severity 25K posts FREE Transfer learning; BERT F1-score: 99%
+Stankevich et al. (2019) Russian depression VKontakte BDI BDI score 531 users UNK Psycholinguistic Markers F1 Score: 65%
 Narynov et al. (2020) Russian depression VKontakte Manual annotation Binary 34K posts FREE
@@
 Ignatiev et al. (2022) Russian depression VKontakte BDI Binary 619 users DUA CatBoost F1 Score: 69%
-Rathnayake and Arachchige (2021) Sinhala depression Twitter, Facebook Manual annotation Binary 1K posts UNK KNN Accuracy: 70%
+Rathnayake and Arachchige (2021) Sinhala depression Twitter, Facebook Manual annotation Binary 1K posts UNK KNN F1-score: 68.5%
 EmoMent Atapattu et al.
 (2022) Sinhala, English mental illness Facebook Manual annotation mental illness, sadness, suicidal, anxiety/stress, psychosomatic, other, irrelevant 2.8K posts AUTH RoBERTa F1 Score: 76%
-Herath and Wijayasiriwardhane (2024) Sinhala suicide Facebook Manual annotation Binary 300 posts UNK Naive Bayes Accuracy: 79%
+Herath and Wijayasiriwardhane (2024) Sinhala suicidal ideation Facebook Manual annotation Binary 300 posts UNK SVM F1-score: 76%
 Leis et al. (2019) Spanish depression Twitter Self-disclosure, manual annotation Binary 540 users, 1K posts FREE
-SAD López-Úbeda et al. (2019) Spanish anorexia Twitter Hashtags Binary 5.7K posts FREE SVM Accuracy: 91.6%
-Valeriano et al. (2020) Spanish suicide Twitter Manual annotation Binary 2K posts FREE Word2Vec; LR Accuracy: 79%
-Ramírez-Cifuentes et al. (2020) Spanish suicide Twitter Manual annotation Binary 252 users N/A
-Ramírez-Cifuentes et al. (2021) Spanish anorexia Twitter Manual annotation Anorexia, control, under treatment, recovered, doubtful 645 users N/A
+SAD López-Úbeda et al. (2019) Spanish anorexia Twitter Hashtags Binary 5.7K posts FREE SVM F1-score: 91.6%
+Valeriano et al. (2020) Spanish suicidal ideation Twitter Manual annotation Binary 2K posts FREE Word2Vec; LR F1-score: 79%
+Ramírez-Cifuentes et al. (2020) Spanish suicidal ideation Twitter Manual annotation Binary 252 users N/A Psychological features, LR F1-score: 80%
+Ramírez-Cifuentes et al. (2021) Spanish anorexia Twitter Manual annotation Anorexia, control, under treatment, recovered, doubtful 645 users N/A CNN, LR F1-score: 98%
 Villa-Pérez et al.
-(2023) Spanish, English depression, ADHD, anxiety, ASD, bipolar, eating disorders, OCD, PTSD, schizophrenia Twitter Self-disclosure Labels for multiple disorders 6K users DUA N-Grams; XGBoost AUC: 71.2%
+(2023) Spanish, English depression, ADHD, anxiety, ASD, bipolar, eating disorders, OCD, PTSD, schizophrenia Twitter Self-disclosure Labels for multiple disorders 6K users DUA N-Grams; XGBoost F1-score: 82.4%
 MentalRiskES Romero et al.
-(2024) Spanish depression, anxiety, suicide, eating disorders Telegram Manual annotation Binary + suffer + in favour (sf), suffer + against (sa), suffer + other (so) for Depression 1.2K users AUTH Social media text; mDeBERTa F1 Score: 46%
-Cremades et al. (2017) Spanish, English suicide Facebook, Twitter, Blogspot, Reddit, Pinterest Manual annotation Binary 97 posts FREE
-Coello-Guilarte et al. (2019) Spanish, English depression Twitter Self-disclosure Binary 316 users FREE BA-LIWC F1 Score: 65%
+(2024) Spanish depression, anxiety, suicidal ideation, eating disorders Telegram Manual annotation Binary + suffer + in favour (sf), suffer + against (sa), suffer + other (so) for Depression 1.2K users AUTH Social media text; mDeBERTa F1 Score: 46%
+Cremades et al. (2017) Spanish, English suicidal ideation Facebook, Twitter, Blogspot, Reddit, Pinterest Manual annotation Binary 97 posts FREE
+Coello-Guilarte et al. (2019) Spanish, English depression Twitter Self-disclosure Binary 316 users FREE BOW, SVM F1 Score: 78%
 Katchapakirin et al. (2018) Thai depression Facebook TMHQ Binary 35 users UNK RF F1 Score: 88.9%
@@
 Hämäläinen et al. (2021) Thai depression Online blogs Manual annotation Binary 900 posts FREE BERT Accuracy: 77.53%
-Mahasiriakalayot et al. (2022) Thai depression Twitter Manual annotation Depression symptoms 3.1K posts UNK LSTM Accuracy: 89.17%
-Boonyarat et al. (2024) Thai suicide Twitter Manual annotation Binary + 6 emotions 2.4K posts FREE Linguistic features; BERT F1-score: 90%
-Benjachairat et al. (2024) Thai suicide Twitter Manual annotation C-SSRS Labels 5.1K posts AUTH Text features; LSTM F1-score: 93.88%
+Mahasiriakalayot et al. (2022) Thai depression Twitter Manual annotation Depression symptoms 3.1K posts UNK LSTM F1-score: 81.06%
+Boonyarat et al. (2024) Thai suicidal ideation Twitter Manual annotation Binary 2.4K posts FREE Linguistic features; BERT F1-score: 93%
+Benjachairat et al. (2024) Thai suicidal ideation Twitter Manual annotation C-SSRS Labels 5.1K posts AUTH Text features; LSTM F1-score: 93.88%